### v1

# Missing Pieces: How Framing Uncertainty Impacts Longitudinal Trust in AI Decision Aids — A Gig Driver Case Study

Abstract.
Decision aids based on artificial intelligence (AI) are becoming increasingly common. When such systems are deployed in environments with inherent uncertainty, following AI-recommended decisions may lead to a wide range of outcomes. In this work, we investigate how the framing of uncertainty in outcomes impacts users’ longitudinal trust in AI decision aids, which is crucial to ensuring that these systems achieve their intended purposes. More specifically, we use gig driving as a representative domain to address the question: how does exposing uncertainty at different levels of granularity affect the evolution of users’ trust and their willingness to rely on recommended decisions? We report on a longitudinal mixed-methods study where we measured the trust of gig drivers as they interacted with an AI-based schedule recommendation tool. Statistically significant quantitative results indicate that participants’ trust in and willingness to rely on the tool for planning depended on the perceived accuracy of the tool’s estimates; that providing ranged estimates improved trust; and that increasing prediction granularity and using hedging language improved willingness to rely on the tool even when trust was low. Additionally, we report on interviews with participants which revealed a diversity of experiences with the tool, suggesting that AI systems must build trust by going beyond general designs to calibrate the expectations of individual users.
1. Introduction
Typically, decision aids based on artificial intelligence (AI) function by both recommending decisions and estimating how good the outcomes of following those decisions will be. AI decision aids are finding increasing adoption in all aspects of society, not just for organisations in healthcare (Wang et al., 2021), finance (Milana and Ashta, 2021), education (McConvey et al., 2023), commerce (Riahi et al., 2021), criminal justice (Završnik, 2019), and other domains, but also for the communication, expression, and productivity of individuals (Skjuve et al., 2023). Yet, when users follow the decisions of such systems, they can encounter unexpected variability in outcomes, including unintentional harms (McGregor, 2021). This is less likely to be caused by malice than by the uncertainty that inherently results from decision aids’ complex interactions with other people and systems (Russell et al., 2017).
How, then, can we ensure that AI decision aids are able to function in concert with their users to achieve their goal of helping make better decisions? This goal can only be achieved by moving beyond system-centric objectives to also focus on the perceptions and attitudes of end-users (Bansal et al., 2019). In particular, trust has been consistently recognised as a factor that mediates the interactions between AI systems and their users (Lockey et al., 2021). We follow the common definition of trust in AI as “the willingness of a person to be vulnerable to the actions of an AI decision aid, based on the expectation that it will perform a decision-making task important to the trustor” (Mayer et al., 1995; Solberg et al., 2022). It is well-known that the design of AI systems impacts the expectations and thus the trust of users (Burgess et al., 2023; Danner et al., 2020; Jacobs et al., 2021; Kocielnik et al., 2019; Kunkel et al., 2019; Scharowski et al., 2023).
As defined above, trust in an AI decision aid is shaped by context. Specifically, psychological models predict that user have a baseline level of trust in an AI decision aid, the expression of which is modulated by their perception of uncertainty in the outcomes of relying upon the system with some context (Solberg et al., 2022). Thus, the presence of uncertainty is important for ensuring that empirical measurements of users’ trust in AI are ecologically valid. However, for practical or ethical reasons, most empirical evaluations of how trust is impacted by the design of AI decision aids have relied on hypothetical (e.g., medicine) or low-stakes (e.g., video recommendation) deployment contexts (Ueno et al., 2022) devoid of such uncertainty.
In this work, we focus on the deployment context of gig driving, where drivers use their personal vehicles to fulfill ridesourcing and food delivery requests from platforms such as Uber, Lyft, DoorDash, and Instacart. This context combines environmental uncertainty with repeated interactions, as drivers must repeatedly make decisions about which requests to fulfill based on an incomplete view of the platform’s supply and demand. Within this context, we contribute an in situ case study, where we evaluate the effect of different framings of uncertainty on user trust in a AI decision aid. To do so, we deployed a schedule recommendation tool that also provides estimates of potential earnings, and measured gig drivers’ trust in the tool over longitudinal trials. Our goal is to address the following research questions with our findings:
[…]
How do users’ perceptions of an AI decision aid’s accuracy in estimating outcomes impact their trust in it?
[…]
How do users’ perceptions of an AI decision aid’s accuracy in estimating outcomes interact with their trust to affect their reliance on it?
[…]
How does the framing of uncertain predictions impact users’ trust and reliance in an AI decision aid?
2. Related Work
2.1. Trust in AI
Different definitions of trust in AI have been proposed (Gillath et al., 2021; Kaplan et al., 2023; Lee and See, 2004; Mayer et al., 1995; Mcknight et al., 2011). We adopt a general definition based on (Mayer et al., 1995), whose model of organisational trust has been widely adopted by work on human-AI trust (Cai et al., 2019a; Liao and Sundar, 2022b; Meske and Bunde, 2020; Saßmannshausen et al., 2021; Textor et al., 2022). Specifically, we draw from a refinement of their model by (Solberg et al., 2022) that adapts it to the context of AI decision aids. An alternative but also widely-adopted (Aoki, 2021; Nourani et al., 2021; Rader et al., 2018) definition is that of (Lee and See, 2004). Both (Lee and See, 2004; Mayer et al., 1995) distinguish trust (an internal attitude) from reliance (the external, behavioural expression of that attitude). However, (Lee and See, 2004)’s definition is less general, as it entails an expectation of positive outcomes from following an AI system. (Solberg et al., 2022) also distinguishes reliance (allowing the AI to recommend decisions) from compliance (following the AI’s recommended decisions).
2.1.1. Measurement of Trust
Because trust is inherently contextual, no single consistent set of measures has been adopted by the literature. The reviews of (Kohn et al., 2021; Ueno et al., 2022) found a great diversity of measures for trust in automation, but also that experimental design choices were usually not informed by nor validated against models of trust. We help rectify this issue by drawing from (Solberg et al., 2022)’s model of trust in AI decision aids for our study design.
2.1.2. Trust in Challenging Contexts
Past work in human-AI interaction has assessed trust under a variety of contexts (Ueno et al., 2022), with some of them entailing unique challenges that impact trust dynamics. We focus on two such types of challenges in this work.
One challenge that impacts trust in AI is the level of uncertainty, which has also been referred to as risk (Stuck et al., 2021; Solberg et al., 2022). Uncertainty encapsulates different sources of variability that make it difficult for users to reason about the potential outcomes of relying on an AI; specifically, (Stuck et al., 2021) separated uncertainty into situational (i.e., inherent in the decision-making context) and relational (i.e., introduced by the AI) components. Prior studies on relational uncertainty in human-AI interaction have focused on medical (Burgess et al., 2023; Cai et al., 2019a; Beede et al., 2020; Jacobs et al., 2021; Wang et al., 2021; Sivaraman et al., 2023; Yang et al., 2023) and financial (Schoeffer et al., 2022; Scharowski et al., 2023) contexts, and generally relied on simulated data. In parallel to this work, the AI explainability literature has increasingly focused on trust calibration, which aims to give users realistic expectations regarding when and why AI systems may or may not perform well (Bhatt et al., 2021; Tomsett et al., 2020; Zhang et al., 2020). Various experiments have compared the effects of explanation types on trust calibration (Kocielnik et al., 2019; Bansal et al., 2021; Wang and Yin, 2021; Cau et al., 2023; Naiseh et al., 2023).
Another challenge that can dynamically impact trust in AI systems is their repeated interactions with users, the effects of which have received limited consideration in past work on human-AI interaction. (Ueno et al., 2022) reported that most studies in human-AI interaction evaluated measures of trust only once, although some studies have attempted to control temporal variation through randomisation (Danner et al., 2020; Naiseh et al., 2021; Wintersberger et al., 2021). Several studies have accounted for the evolution of trust over time more explicitly by comparing multiple measurements of trust over weeks or months (Jiang et al., 2004; Mou and Cohen, 2017; Mou et al., 2017; Kunkel et al., 2019).
Generally, studies of trust under these challenging contexts have relied on hypothetical or low-stakes scenarios. Even studies of deployed systems (Beede et al., 2020; Wang et al., 2021) have been limited to observational studies that did not compare multiple designs. Within our uncertain, longitudinal, and higher-stakes context, users have varying levels of experience with how environmental uncertainty impacts their earnings, but all of them are freshly experiencing the impacts of the AI’s relational uncertainty. Under this context, we contribute a mixed-methods study that compares the effects of different designs for an AI decision aid on trust. Specifically, we investigate a design dimension not explored by prior studies: how different framings of relational uncertainty in these recommendations using hedging language (Kaufman and Kirsh, 2022) can alter the longitudinal dynamics of trust.
2.2. Recommendation Systems
Recommendation systems are mediators of the interfaces between people and the technologies they use. However, these systems do not always serve to reveal users’ unstated preferences, but may instead influence them over repeated interactions (Gretzel and Fesenmaier, 2006). The dynamics of these interactions are a double-edged sword (Carroll et al., 2022): they may either lead to undesirable concentration (Ferraro et al., 2020; Zhou et al., 2021) or incentivise further exploration of content (Liang and Willemsen, 2022). There has been limited work on how these longitudinal interactions can potentially be shaped by the design of recommendation systems. It has, however, been found in single-shot settings that the specific wording used to frame recommendations can influence their perception by users. (Sundar et al., 2008; Liao and Sundar, 2022a) found that framing recommendations in terms of similarity to other users improves perceived quality and trust (i.e., the bandwagon effect). (Komiak and Benbasat, 2006; Sun et al., 2023; Suwanaposee et al., 2023) recovered conflicting conclusions on whether personalised framings of recommendations improve perceived quality (i.e., the Barnum effect). Our investigation of hedging addresses the question of how such effects can interact with users’ experiences to impact trust in recommendations longitudinally.
2.3. Gig Work and Gig Driving
Platform-based gig work is among the most prevalent contexts today in which AI-based systems make managerial decisions that control individuals (Holmström and Hällgren, 2022). Gig work offers workers the flexibility and autonomy to choose when and where they want to work (Lee et al., 2015). However, a growing body of work has demonstrated that this autonomy is hampered by the opacity of platforms’ assignment, pricing, and evaluation mechanisms (Griesbach et al., 2019; Lee et al., 2015; Singh et al., 2023; Tan et al., 2021; Wood et al., 2019). We follow a line of work that has responded to this issue by envisioning novel, worker-oriented designs of platform infrastructures (Stein et al., 2023; Zhang et al., 2022).
In particular, we tackle the challenge of information asymmetry in pricing mechanisms for gig driving platforms (Maffie, 2023; Rosenblat and Stark, 2016), which dynamically vary compensation based on consumer demand (Lee et al., 2015) and gamification (Zhang et al., 2022). Platforms have given drivers little visibility into the demand patterns and pricing structures that dictate their earnings (Griesbach et al., 2019; Lee et al., 2015), causing difficulties in planning and inequities in revenue (Miao et al., 2023). The closest-related works to our own are those of (Battifarano and Qian, 2019; Khan et al., 2022; Zhang et al., 2023), who designed tools that provide transparency into earnings. We go beyond these works by deploying such a tool as part of an AI decision aid that also gives schedule recommendations.
3. Schedule Recommendation Tool
In this section, we develop an AI-based schedule recommendation tool for gig drivers. To our knowledge, such a schedule recommendation tool has not been proposed in prior AI literature. We use this tool as an example AI decision aid for our study on the relationship between the framing of uncertainty in outcomes and the longitudinal trust in AI decision aids.
3.1. Decision Aid Design
We focus on one aspect of activity planning for gig drivers: choosing when to work. Most drivers seek to earn as much as they can from driving. However, the importance of this goal varies; some drivers rely on gig driving as a primary or sole source of income, while others treat it as a source of supplemental income (Ma and Hanrahan, 2019; Wu et al., 2019). The amount of time that drivers invest also varies; part-time drivers balance gig driving with other responsibilities, while others view it as a full-time commitment (Berger et al., 2020; Lee et al., 2015; Ma and Hanrahan, 2019). Our schedule recommendation tool is designed to help gig drivers decide when to work given their constraints and preferences. The tool consists of two modules.
First, an estimation module prospectively predicts , the earnings that drivers can expect during a specific hour on a specific weekday . These estimates could be computed by a machine learning model or directly averaged from historical data.
Second, a scheduling module takes the estimated earnings and each driver’s constraints as inputs to produce an optimal set of working times that leads to the highest possible earnings. This module solves a constrained optimisation problem:
where denotes hour on weekday , denotes whether time slot is recommended, is the expected earning from driving during slot from the estimation module, is a binary indicator of whether the driver is available during slot , is the maximum number of hours they wish to work on weekday , and is the maximum number of hours they wish to work per week. Following past work on gig driving (Griesbach et al., 2019; Lee et al., 2015; Maffie, 2023; Zhang et al., 2023), we focus on daily and weekly constraints for driving hours and earnings.
Some drivers who value earnings to a greater extent set personal targets for their hourly or daily earnings (Zhang et al., 2022). In this case, we propose to solve the following alternative problem:
where is the minimum amount the driver wishes to earn on weekday , and is the minimum amount the driver wishes to earn during the entire week.
3.2. Interface Design
Next, since our proposed AI decision aid needs to take into account each driver’s constraints, we sought to design a front-end interface that would allow drivers both to enter these constraints and to view the optimised schedule of working times generated by the tool based on them. Our design consists of a two-page interface.
First, a constraint page (Figure S1 in Appendix D) elicits the constraints needed for Equation 1 or Equation 2 by prompting users to enter their availability and goals. The page first asks users to select an optimisation objective: whether they would like set “a maximum number of hours to drive” ( in Equation 1) or “a minimum amount of money to make” ( in Equation 2). Then, it asks users to choose to set daily ( and ) or weekly ( and ) constraints. To maximise users’ perceived control over the system (see Section 4.1), we allowed them to choose these options freely as opposed to enforcing them within conditions. Lastly, the page elicits availability information for each hour of each day ( in Equations 1 and 2) in a tabular format. Unlike (Zhang et al., 2023), we account for the fact that users’ availability may change between days.
Second, a schedule page (Figure 1) presents a tabular schedule generated by solving Equation 1 or Equation 2 to the user. The table shows the estimated earning of each time slot in its corresponding cell; the solution is shown by highlighting the recommended time slots (i.e., the ones that lead to the highest earnings subject to constraints). A summary of the user’s expected earnings for the week is shown above the table. In the interests of maximising perceived control, we allowed users to revisit the constraint page until they were satisfied with the generated schedule.
4. Uncertainty and Trust
4.1. Research Hypotheses
Provided that all parameters are specified, Equations 1 and 2 can be easily solved to generate a recommended schedule. However, this problem entails environmental uncertainty: each time slot’s earnings will dynamically vary with demand and supply. Inaccurate earnings estimates could substantially impact the quality of the solution (Cameron et al., 2022). Furthermore, information asymmetry prevents drivers and our AI decision aid from observing the ground truth distribution of gigs in real time. Thus, further uncertainty is introduced by the AI system having to rely on historical data or other proxies. How does this uncertainty impact trust and reliance?
Regarding Research Questions 1 and 2, (Solberg et al., 2022)’s model proposes that the outcomes of reliance on an AI decision aid can impact users in two ways. First, they may affect perceptions of the system’s fundamental trustworthiness, in terms of performance (i.e., competency in tasks), process (i.e., transparency of behaviour), and purpose (helpfulness for beneficiaries) (Lee and Moray, 1992). Second, they may affect perceptions of the uncertainty present in a given context (situational risk), which modulates the expression of trust through reliance. Even if the user trusts an AI decision aid in general, they may still feel that trusting it in particularly uncertain contexts could lead to adverse outcomes that fall short of the AI’s promises. If the tool is able to accurately estimate the outcomes of users’ reliance, it may both enhance performance-based trust and lead to higher reliance in uncertain contexts. This leads us to the following two hypotheses:
[…]
Users who perceived the outcomes as meeting or exceeding estimates will have a greater level of performance-based trust than those who perceived the outcomes as falling below estimates.
[…]
Users with the same level of trust prior to interaction will exhibit higher reliance if they have perceived the outcomes as meeting or exceeding estimates.
Meanwhile, regarding Research Question 3, (Solberg et al., 2022)’s model predicts that the design of an AI decision aid impacts its trustworthiness. Which design dimensions are most likely to affect trust? (Holmström and Hällgren, 2022) identified two axes of AI system designs that define different types of human-AI interaction: algorithmic management, the extent of the AI’s control on tracking, evaluation, and decision-making; and transparency, the extent to which these processes can be understood. Along the axis of algorithmic management, (Solberg et al., 2022) hypothesised that high levels of perceived control over otherwise untrustworthy systems may lead to higher levels of trust (Ho et al., 2013). Along the axis of transparency, (Solberg et al., 2022) considered explainability methods as enhancing process-based trust, in accordance with empirical results (Ferrario and Loi, 2022; Shin, 2021), and increased visibility into error rates as harming performance-based trust. We believe that this latter design dimension modulates how users’ perceptions of uncertainty impact their trust and reliance. This leads us to the following two hypotheses:
[…]
Designs which present estimates with greater granularity but not error visibility increase users’ trust and reliance.
[…]
Designs which show ranged estimates decrease users’ trust and reliance through increasing error visibility.
4.2. Tool Conditions
Screenshot of a box titled “Shift Recommendation Tool” on the schedule page, showing Condition 1. An estimated weekly earning of $1 412 is highlighted with box A. Inset A shows variants of how the weekly earning is presented (Condition 2, a table showing “An average of…” with $210 for Monday, $194 for Tuesday, etc.; Conditions 3 and 4, a table showing “On an average week: $1 412”, “On a bad week: $391”, and “On a good week: $3 032”). A recommended schedule shows average earnings for each weekday and each hour. Two cells reading “Average: $15” and “Average: $12” are highlighted with box B. Inset B shows a variant of how the hourly earnings are presented (Conditions 3 and 4, cells reading “Average: $15, Low: $4, High: $34” and “Average: $12, Low: $4, High: $28”). Before the weekly estimate of $1 412, box C encloses text reading “Based on historical data, it is estimated that you will earn a weekly average of”. Inset C shows a variant of this text (Condition 4: “In general, every driver may encounter different conditions while they are driving, which leads to differences in earnings from individual to individual and from one day to the next. On average, based on historical data, a driver following this schedule will earn”).
In service of Research Question 3, we varied the design of the schedule page in our tool between 4 different framings of uncertainty. In 2 of the 4 designs, we accomplished this by presenting earning estimates as ranges instead of as a point estimate. The range consisted of 3 quantities: a mean earning ( for hourly earnings; for weekly earnings), a pessimistic earning (based on the 10th percentile of earnings), and an optimistic earning (based on the 90th percentile of earnings). Figure 1 shows the 4 conditions:
Reference condition. Participants were only shown their mean weekly estimated earnings and their mean estimated earnings for each hour during the week.
Daily estimates. To assess the effect of increasing granularity but not error rate visibility, participants were shown their mean estimated earning for each day instead of their mean estimated weekly earning. The schedule still showed mean estimated earnings for each hour during the week.
Ranged estimates. To assess the effect of increasing error rate visibility, participants were shown mean, pessimistic, and optimistic estimated earnings for the entire week, as well as for each hour during the week.
Ranged and hedged estimates. To assess the effect of hedging on error rate visibility, the textual description of the estimates was changed from Condition 3. Instead of “Based on historical data, it is estimated that you will earn”, Condition 4 states “On average, based on historical data, a driver following this schedule will earn”.
5. Pilot Interviews
To assess the utility of the AI gig recommendation tool design introduced in Section 3, we created a prototype based on them. We then conducted a series of interviews to understand gig drivers’ needs and how well the prototype aligned with them. The interview methodology was approved by our Institutional Review Board (IRB).
5.1. Tool Design
Following common practice in UX design (Feng et al., 2023), we used Figma (Figma, Inc., nd) to create the prototype design (Appendix E).
The constraint page’s design (Figure S6) was largely similar to the final design. However, the prototype had a monotone colour scheme and had instructions that were worded less clearly, which were improved based on feedback from the pilot (Section 5.3). Additionally, due to technical limitations, dropdowns were used in place of text boxes, and no input validation was present. However, we consider the impact of these limitations to be minor given that the interviewer controlled the page.
The schedule page’s design (Figure S7) was kept static for the pilot to gather more uniform feedback from participants. Like the constraint page, the prototype differed from the final design in its colour scheme and clarity of wording. The prototype was most similar to Condition 3 (Section 4.2) in that it displayed a range consisting of mean (“On an average week…”), pessimistic (“On a bad week…”), and optimistic (“On a good week…”) weekly earnings above the tabular schedule. Unlike the final design of Condition 3, however, ranges were not shown for hourly estimates; this was added based on feedback from the pilot (Section 5.3).
5.2. Methodology
Participants began by completing a web-based consent form and a demographics survey that collected their age, gender, and education level. This form was separate from the one for the longitudinal trials (Section 6.1.1). After completing the web-based portion, participants were invited via email to complete 20–30-minute audio-recorded Zoom interviews conducted by a single author. Participants were compensated with a $10 Amazon gift card.
In the first 5–10 minutes, the interview focused on formatively understanding drivers’ needs and motivations. We asked the participants a series of questions about their needs and motivations, including what they liked or disliked about driving for gig platforms, their typical goals and routines when driving, and the extent to which they would find schedule recommendations useful, along with demographic questions.
In the last 15–20 minutes, the interview focused on evaluatively understanding how well the tool met drivers’ needs. To ensure a consistent experience, the Figma prototype was opened in the interviewer’s browser and shown to participants in a screensharing session. First, on the constraint page, the participant was asked to work with the interviewer to interact with the page, entering the constraints as if they were using the tool for their actual planning. Then, on the schedule page, the participant was shown a schedule with mocked earning estimates. Finally, the participant was asked about their overall opinions of the tool. The full interview script is shown in Section B.1.
Transcripts for the interviews were automatically generated by Zoom. The author who conducted the interviews first reviewed and corrected these transcripts, then performed structural coding (Saldaña, 2015a). Afterwards, two authors separately used QualCoder 3.3 (Curtain, 2023) to perform open coding (through descriptive, process, and in vivo coding (Saldaña, 2015a)) and axial coding (Saldaña, 2015b). The two authors then met to reconcile their codes and construct a unified codebook. Finally, the interviewing author re-applied the updated set of codes. Following (Cai et al., 2019b), we did not compute inter-rater reliability, since our grounded analysis used codes as an intermediary to generate theory.
5.2.1. Participants
As with the longitudinal trials (Section 6.1.3), participants were recruited from the user base of Gridwise, a mobile assistant app for gig drivers, in July 2023. We chose this user base instead of recruiting from a particular platform (Zhang et al., 2023) or from social media (Zhang et al., 2022), so as to capture the needs of a diverse set of drivers. Gridwise distributed recruitment messages to 500 users, but otherwise did not interact with participants. Recipients were sampled from users in the United States who had completed at least one gig in a platform linked to the Gridwise app over the week preceding recruitment. We recruited 4 interview participants, hereafter designated as P1–P4:
P1: A 39-year-old female with a professional degree who drives exclusively for delivery platforms
P2: A 55-year-old male with less than a high school degree who drives exclusively for delivery platforms
P3: A 29-year-old female with an undergraduate degree who drives more frequently for ridesharing platforms
P4: A 53-year-old male with a professional degree who drives more frequently for delivery platforms
5.3. Results
The four participants reported a diversity of motivations and routines. While all four participants had specific earning goals, P1, P2, and P4 considered their goals to be important and valued the time flexibility of gig work, whereas P3 was more motivated by the opportunity for human interaction. P1, P2 and P3 had typical times that they drive at, but P1, P2, and P4 also adjusted their driving schedules based on gig availability. All four participants had encountered difficulties in planning due to the unpredictability of rider demand and/or driver supply (with P1, P2, and P3 feeling that gig platforms provide insufficient information), and indicated that they would find schedule recommendations to be useful.
All four participants found the initial design of the tool to be generally understandable, and felt that it would be generally useful for drivers in planning. P1 and P2 liked the fact that the tool presents information to them in a way that reduces the need for guesswork while driving. In particular, P1 suggested that the tool would help mitigate a catch-22: it is not possible to view gig demand information in DoorDash without exiting their Dash (scheduled work period), but doing so seemingly deprioritises them.
On the constraint page, P1 and P4 indicated that the questions aligned well with their goals. P1 and P3 suggested that they would not set the constraints to perfectly align with their routines, so as to receive more information from the tool. Meanwhile, on the schedule page, all four participants liked the estimated hourly earnings, with P1, P2, and P3 indicating that they would be helpful in deciding whether or not to work at particular times of day. Yet, P2, P3, and P4 acknowledged that the estimates were only guesses. P1 and P4 also liked the range of weekly earnings, but P3 felt it assumed they would follow the recommended schedule perfectly. P2 and P3 mentioned that ranges for hourly earnings would be useful.
P1, P2, and P3 all felt that it was better for the tool to have a simple, easy-to-use design. All three indicated that the prototype fulfilled this requirement, although P3 suggested that the wording and design improvements would be necessary (in particular, they felt that the monotone colour scheme of the tool was confusing). P2 and P3 felt that the design needed to be mobile-friendly. The participants also mentioned other desiderata:
More granular constraints. P1, P3, and P4 all suggested ways to limit the scope of the historical gigs used to estimate their earnings. P3 wanted the tool to clarify that the historical data was limited to the city they drive in, and also how recent the data was. P1 and P3 wanted to limit the maximum distance of the historical gigs from their starting point. P4, who works for less popular delivery platforms, wanted to select which platforms the historical gigs came from, and indicated that this would improve their control.
Feedback on performance. P1 and P2 both wanted a way to compare the tool’s estimates with their actual earnings. Regardless of how the estimates compared to reality, P1 suggested that they would be motivating; both P1 and P2 remarked that they have gamified their gig-driving experiences to compare against either themselves or others. P2 also wanted to compare the estimates with their expenses, and P3 wanted a way to view the overall supply of drivers.
6. Longitudinal Trials
After our pilot study, through which we verified the problem formulation and improved the design of our AI decision aid (Section 5), we sought to deploy the decision aid to a larger sample of users. Our goal was to measure how their trust in the tool would evolve over time and vary with different framings of uncertainty. To this end, we conducted a longitudinal trial in which participants interacted with the tool for 7 days over a 14-day period, with the longer time window meant to accommodate variability in participants’ availability. The methodology for this longitudinal trial was approved by our Institutional Review Board (IRB).
6.1. Methodology
6.1.1. Trial Design
The trial was administered through a custom website built using Django 4.1 (Django, 2023), which was designed to be functional on both computers and mobile devices. Crucially, the website was designed to be visually generic and distinct from any apps or websites that participants may use to access gig platforms; this was done to avoid the possibility that participants’ experiences with any individual platform could have biased their perception of the tool. Prospective participants received a link to the website from the recruitment message, which led them to complete the consent form and begin the trial.
On Day 0 of the trial, participants first completed the 12-question Survey 1 — Formative Questions (LABEL:{sec:app:survey:1}). These questions were based on our formative interviews about participants’ needs and motivations from the pilot interviews (Section 5.2). To filter the data used to generate estimates, we also asked participants about the region they drive in and the platforms they work for.
Next, they were directed to a page on which they interacted with the schedule recommendation tool. To mitigate response bias (Dell et al., 2012; Ko et al., 2015), we avoided referring to the tool as having been designed by us; isolated the tool in an iframe; and differentiated the fonts of the website and tool. First, participants filled in their availability and goals on the constraint page. Second, on the schedule page, we assigned participants to one of the 4 conditions to view their schedule. Assignment was done such that each condition had an approximately equal number of participants.
The estimates were generated using gig data from August 2023 in four cities: Los Angeles, New York, Chicago, and Houston. For each city, our data included approximately 100 000–300 000 samples distributed evenly across times and weekdays. Earnings for each time slot (weekday , hour ) were estimated by the mean of what drivers historically earned in this slot, filtered to the participant’s city and platforms. We used this simple, static estimator to focus on the effects of framing uncertainty in the estimates. Our results are intended, however, to generalise to any possible estimator for , including those that generate dynamic estimates based on machine learning. To this end, we intentionally chose not to emphasise to the participants that the schedule page was generated using AI or optimisation, and did not provide any additional information about the data used to generate the estimates.
We sought to measure participants’ trust in the tool throughout the longitudinal trial. As reviewed by (Kohn et al., 2021), internal trust is typically measured by self-report measures. Two widely-used instruments for self-reported trust are (Madsen and Gregor, 2000)’s Human-Computer Trust Questionnaire and (Jian et al., 2000)’s Trust in Automation Scale. The former is a 25-item scale that measures 5 facets of trust (Perceived Understandability, Perceived Technical Competence, Perceived Reliability, Personal Attachment, and Faith) using 5 questions each. Meanwhile, the latter is a 12-item scale that measures both trust and distrust. Following the recommendations of (Kohn et al., 2021), we adapted these validated instruments to our specific context instead of creating custom measures of trust (see comments in Appendix A). To reduce user burden, we considered it infeasible to repeatedly administer these instruments in their entirety, and instead extracted subsets of them for our surveys as recommended by (Kohn et al., 2021). The first measurement of trust took place immediately after participants saw the recommended schedule on Day 0. Survey 2 — Pre-Interaction Trust (Section A.2) consisted of 5 items, one taken from each of (Madsen and Gregor, 2000)’s 5 facets of trust.
After completing Survey 2, participants began their 7 days of interaction with the tool starting on their next available date. On each of Day 1 to Day 7, participants first received their recommended schedule for that particular day, sent via an email scheduled for 30 minutes before the start of their indicated availability. This was intended to present participants with information from the tool right as they were deciding their driving schedule. During the day, participants continued uninterrupted with their driving activity; we repeatedly emphasised that compliance with the recommended schedule was not a condition of full participation. Finally, at the end of their indicated availability, a second scheduled email sent them a link to Survey 3 — Daily Trust (Section A.3). This was an 8-question survey that included items extracted from 3 of (Madsen and Gregor, 2000)’s 5 facets of trust, along with questions on how the participant’s earnings compared to their expectations and the tool’s estimates, and to what extent they intended to rely on the tool for the following day.
If the participant indicated that they intended to continue relying on the tool, they were then presented with a daily variant of the schedule page, showing the subset of the recommended schedule corresponding to the following day. We also allowed them to alter their availability for the coming day using a daily variant of the constraint page. Updated schedules were generated by fixing the recommended time slots for previous days using equality constraints and then re-solving Equation 1 or Equation 2. On the other hand, if the participant’s answer indicated that they intended to pause their interaction for one day, a scheduled email was sent on the following day, which prompted them to either review the next day’s schedule or to pause for an additional day.
On Day 7, we removed the last question of Survey 3 and added Survey 4 — Post-Interaction Trust (Section A.4), a series of 10 questions intended as a retrospective measurement of trust after participants had completed interaction with the tool. This survey included 5 questions drawn from (Jian et al., 2000)’s instrument. We chose not to include questions from this instrument earlier in the trial, as we felt that questions measuring distrust could have biased participants’ reliance. After completing Survey 4, participants were sent a final email that invited them to participate in the interview.
Based on a pilot of the longitudinal trial conducted with 7 participants in August 2023, we determined that the sequence of surveys on Day 0 took an average of 14 minutes and 27 seconds (including idle time), Survey 3 took an average of 2 minutes and 23 seconds per day, and Survey 4 took an average of 2 minutes and 21 seconds. The results from this pilot were excluded from analysis, and the 7 participants were excluded from future recruitment. These figures led us to set the compensation as an $40 Amazon gift card for full study completion: $6 for the Day 0 surveys, $2 for each daily survey, and a $20 completion bonus. For this payment structure, we drew from (van Berkel and Kostakos, 2021)’s recommendation of micro-compensation. However, we made one payment upon either study completion or the passage of 14 days to mitigate effects on ecological validity.
6.1.2. Interview Design
For participants who indicated their desire to be interviewed, an audio-recorded Zoom interview of 20–30 minutes was conducted by a single author. The interview was not a condition of full study completion. Participants were compensated separately with a $20 Amazon gift card.
Overall, the interview focused on assessing dimensions of participants’ experiences that were not evident from the surveys, such as the thought processes underlying their perceptions of the tool and their decisions regarding reliance and compliance. The interview began with questions about participants’ motivations and routines, which led into questions assessing the constraint page’s alignment with their decision-making process as well as their perceived control over the schedule through this page. Next, participants were asked about their understanding of the schedules, the extent of their reliance on and compliance with them, their perceptions of the outcomes, and whether these changed throughout their 7 days of interaction with the tool. Another set of questions then focused on the estimates, including perceptions of how accurate they were and whether participants would have preferred a different condition. We also asked participants to recall a specific day of interaction with the tool, in terms of how the schedule and estimates affected their behaviour for that day and for the following day. Finally, participants were asked for their overall thoughts on the tool.
6.1.3. Participants
Participants were recruited from the user base of Gridwise in September 2023. We chose to recruit from this user base again to access a larger and more diverse sample of both historical data and participants. Gridwise distributed recruitment messages to users who had completed at least one gig in DoorDash, Grubhub, Instacart, Lyft, Uber, or Uber Eats (the six platforms for which historical data was available) over the month preceding recruitment, and resided in one of the the four cities with the historical data used to generate the tool’s earning estimates (see Section 6.1.1): Los Angeles, New York, Chicago, and Houston.
Bar graph showing the number of participants from Day 0 to Day 7, followed by the Interview. Day 0 had Survey 1, the constraint and schedule pages, and Survey 2 as associated activities. On Day 0, the four conditions respectively had 13, 14, 12, and 12 participants. Days 1–7 each had the daily schedule and Survey 3 as associated activities; Day 7 also had Survey 4. Over the following week, retention for Conditions 2 and 4 remained high, but retention dropped substantially for Conditions 1 and 3. By Day 7, the four conditions respectively had 6, 10, 6, and 11 participants. The Interview had 0, 3, 1, and 3 participants for the four conditions.
Of the recipients, 51 completed at least the sequence of surveys on Day 0 and thus qualified for compensation. Among them, 25 (49%) were from Los Angeles, 10 (19.6%) were from New York, 8 (15.7%) were from Chicago, and 8 (15.7%) were from Houston; 4 (7.8%) were aged 18–24, 15 (29.4%) were aged 25–34, 22 (43.13%) were aged 35–44, 8 (15.7%) were aged 45–54, and 2 (3.9%) were aged over 55; 34 (66.7%) were male, 15 (29.4%) were female, and 1 (2%) was non-binary; 5 (9.8%) had a graduate degree, another 16 (31.4%) had an undergraduate degree, another 11 (21.6%) had a professional degree, and another 19 (37.3%) had a high school degree. We did not collect race and ethnicity data, as this is a rare practice that is not directly relevant to our research questions (Chen et al., 2023).
Among these, 7 of these completed the interview and received an additional $20 of compensation. They were:
F1: A 47-year-old male with a high school degree who drives for Instacart and Lyft in Los Angeles, Condition 2
F2: A 42-year-old female with a high school degree who drives for Uber Eats in Los Angeles, Condition 2
F3: A 29-year-old female with a graduate degree who drives for Lyft in Chicago, Condition 4
F4: A 49-year-old male with an undergraduate degree who drives for Lyft and Uber in Los Angeles, Condition 4
F5: A 46-year-old male with an undergraduate degree who drives for DoorDash, GrubHub, and Uber Eats in Chicago, Condition 3
F6: A 39-year-old male with a graduate degree who drives for Lyft, Uber, and Uber Eats in Los Angeles, Condition 2
F7: A 39-year-old male with a professional degree who delivers for DoorDash in Los Angeles by bike, Condition 4
6.1.4. Analysis
For quantitative analysis, we applied statistical models implemented in statsmodels 0.14.0 (Seabold and Perktold, 2010) to our measurements. The questions comprising the instruments were presented to participants as 5-point Likert-type scales in ascending order from 1–5, following (Chyung et al., 2017, 2018). Ordinary least squares (OLS) models were fitted to pre- and post-interaction (Survey 2 and Survey 4) trust scores. For Survey 2 trust scores, we took the average of the 5 items drawn from (Madsen and Gregor, 2000); for the Survey 4 trust score, we first inverted 2 items measuring distrust from (Jian et al., 2000) and then took their average with 3 other items from (Jian et al., 2000). To model intra-subject dependencies, linear mixed models (LMMs) were fitted to daily data from Survey 3, with participant IDs as random effects. For the trust score, we took the average of the 3 items drawn from (Madsen and Gregor, 2000). We also computed a reliance indicator, which was 0 if a participant (a) indicated their desire to decrease or pause their use of the tool on the following day or (b) failed to complete the following day’s Survey 3, and 1 otherwise. For each model, we started with all relevant variables, and then manually performed variable selection to minimise the Akaike information criterion (AIC) (Cavanaugh and Neath, 2019). The reliance indicator is more properly fit by a logistic generalised linear mixed model (GLMM). However, since the statsmodels implementation of GLMMs did not support either -values or the AIC, model selection was performed using the LMM equivalent. We report results for both.
We used the same methodology as the pilot interviews (Section 5.2) to analyse the interviews for the longitudinal trials.
6.2. Results
Of the 51 participants, 33 went on to complete all 7 days of interaction with the tool, thus receiving $40 of compensation. We show full retention statistics for each of the 4 conditions in Figure 2. Overall, although individual participants differed in their level of reliance on the recommendations (see also Section 6.2.2), participants generally felt that our estimates were reasonably calibrated. Across all daily interactions corresponding to Survey 3 responses, participants reported that their earnings for the day felt roughly equal to the tool’s estimate after 49% of interactions, higher than the tool’s estimate after 23% of interactions, and lower than the tool’s estimate after 28% of interactions — a roughly symmetric distribution. Nevertheless, trust and reliance still significantly differed between conditions, as is also evident from the differences in retention rates in Figure 2. We now explore these differences through our quantitative results.
6.2.1. Quantitative Results
In describing the statistical models that we fitted, we use the following variable notations: for the Survey 2 trust score; for the Survey 3 trust score on day ; for the Survey 4 trust score; for the reliance indicator on day ; for the participant’s schedule page condition; for the monetary difference between a participant’s actual and estimated earnings for the hours that they worked on day ; for the monetary difference between a participant’s actual and estimated earnings for the hours that the tool recommended on day ; for an indicator denoting whether the participant perceived their earnings to be at least as high as the tool’s estimate; for an indicator denoting whether the participant perceived their earnings to be at least as high as their expectations; and as the day for a data point from Survey 3. Statistically significant coefficients are denoted using asterisks for each significance level: * (0.05), ** (0.01), *** (0.001).
Research Question 1
This research question assesses the relationship between users’ perceived accuracy of the tool’s estimates and trust. As applied to our context, Hypothesis 1 hypothesises that trust would be increased by users perceiving their earnings as being at least as good as the tool’s estimates. Based on the trust scores self-reported during and after tool interactions (Surveys 3 and 4), we consider Hypothesis 1 to be partially supported.
For the daily Survey 3 trust score, the best-fitting LMM had as fixed effects, and thus depended only on interactions of the day with (1) the monetary difference in earnings for the participant’s worked hours and (2) the perceived accuracy of the tool’s estimate. When participants perceived their earnings to be at least as high as the tool’s estimates, their trust scores were significantly higher (; vs ). We also observed significant longitudinal effects: the trust score generally increased with each passing day, but the rate of increase was significantly lessened among daily interactions where participants perceived the tool’s estimates as being accurate (; vs ). This result can likely be attributed to the higher baseline level of trust for participants who perceived the tool’s estimates as being accurate. Fitted coefficients are shown in Table 1.
Coefficient | |||
Fixed effects | |||
Intercept | 3.158 | 0.233 | 0.000*** |
0.133 | 0.044 | 0.003** | |
0.004 | 0.002 | 0.004** | |
-0.001 | 0.000 | 0.010** | |
0.543 | 0.233 | 0.020* | |
-0.122 | 0.050 | 0.014* | |
Random effect | 0.527 | 0.258 |
Table of linear mixed model coefficients (intercept, fixed effects, and random effects) for daily Survey 3 trust. Significant coefficients are the intercept (3.158), (0.133), (0.004), the interaction of and (-0.001), (0.543), and the interaction of and (-0.122).
However, we failed to find significant evidence that the outcomes of individual interactions improved participants’ trust when considering their experiences retrospectively. For the post-interaction Survey 4 trust score, the best-fitting OLS model was simply , and thus only depended on the perceived accuracy of the tool’s estimate for the first day. This variable did not have a significant effect (; vs ). Still, the positive correlation of this variable with post-interaction trust echoes the findings from the longitudinal studies of (Mou and Cohen, 2017; Mou et al., 2017) regarding the importance of initial experiences. Fitted coefficients are shown in Table 2.
Coefficient | |||
Intercept | 3.3600 | 0.292 | 0.000*** |
0.4900 | 0.317 | 0.132 |
Table of ordinary least squares model coefficients for post-interaction Survey 4 trust. The intercept (3.36) is significant, but the coefficient for (0.49) is not.
Research Question 2
This research question assesses the relationship between users’ perceived accuracy of the tool’s estimates, trust, and reliance. As applied to our context, Hypothesis 2 hypothesises that, holding trust constant, reliance would be increased by users perceiving their earnings as being at least as good as the tool’s estimate. Based on the daily Survey 3, we consider Hypothesis 2 to be supported. For the Survey 3 reliance indicator (whether participants increased or maintained their intended level of reliance), the best-fitting LMM that included the trust score had as fixed effects, and thus only depended on (1) the perceived accuracy of the tool’s estimate, (2) the perceived accuracy of the participant’s expectations, and (3) the trust score. Removing the trust score improved the AIC; however, we retained it to isolate the effect of perceived accuracy on reliance independent of trust. Holding the trust score constant, we found significant evidence that reliance was higher when participants perceived the tool’s estimates as being accurate (; LMM means vs ; GLMM log-odds vs ). Fitted coefficients for both the LMM and GLMM are shown in Table 3. However, -values for the GLMM could not be computed.
LMM | GLMM | ||||
Coefficient | |||||
Fixed effects | |||||
Intercept | 0.411 | 0.140 | 0.003** | -1.140 | 0.6619 |
0.214 | 0.073 | 0.003** | 1.530 | 0.457 | |
0.069 | 0.066 | 0.299 | -0.066 | 0.465 | |
0.029 | 0.033 | 0.389 | 0.412 | 0.178 | |
Random effect | 0.074 | 0.088 | -43.000 | 0.181 |
Table of linear mixed model and generalised linear mixed model coefficients (intercept, fixed effects, and random effects) for daily Survey 3 reliance. Significant coefficients are the intercept (0.411 for the LMM, -1.14 for the GLMM) and (0.214 for the LMM, 1.53 for the GLMM), where significance is based on the LMM -values only.
Research Question 3
This research question assesses the effect of framings of uncertainty on trust and reliance by comparing them across the different conditions outlined in Section 4.2. Relative to the baseline of Condition 1, Hypothesis 3 hypothesises that designs which present estimates with higher granularity but not error visibility (Condition 2) increase trust and reliance, while Hypothesis 4 hypothesises that designs which present ranges for estimates (Conditions 3 and 4) decrease trust and reliance. Based on the trust scores and reliance indicators from all surveys, we consider Hypothesis 3 and Hypothesis 4 to be partially supported.
For the pre-interaction Survey 2 trust score, the best-fitting OLS model was , and thus depended only on the condition. Group means and pairwise comparisons are shown in Table 4. Pairwise comparison means are expressed as the mean of the second group minus the mean of the first.
Coefficient | |||
Group means | |||
Condition 1 | 3.339 | 0.266 | 0.000*** |
Condition 2 | 3.629 | 0.369 | 0.000*** |
Condition 3 | 4.183 | 0.384 | 0.000*** |
Condition 4 | 3.367 | 0.384 | 0.000*** |
Pairwise comparisons | |||
1 - 2 | 0.290 | 0.369 | 0.432 |
1 - 3 | 0.845 | 0.384 | 0.028* |
1 - 4 | 0.028 | 0.384 | 0.941 |
2 - 3 | 0.555 | 0.377 | 0.141 |
2 - 4 | -0.262 | 0.377 | 0.487 |
3 - 4 | -0.817 | 0.391 | 0.037* |
Table of pairwise comparisons for the ordinary least squares model for pre-interaction Survey 2 trust. The group means for the four conditions are 3.339, 3.629, 4.183, and 3.367. Significant differences exist in Conditions 1 and 3 (difference 0.845) and Conditions 3 and 4 (difference -0.817).
For the daily Survey 3 trust score, the best-fitting LMM including the condition had as fixed effects, and thus depended only on (1) the condition and interactions of the day with (2) the monetary difference in earnings for the participant’s worked hours and (3) the perceived accuracy of the tool’s estimate. Note the lack of interactions for the condition. Group means, pairwise comparisons, and other coefficients are shown in Table 5.
Coefficient | |||
Group means | |||
Condition 1 | 3.160 | 0.316 | 0.000*** |
Condition 2 | 3.083 | 0.378 | 0.000*** |
Condition 3 | 3.520 | 0.385 | 0.000*** |
Condition 4 | 2.770 | 0.381 | 0.000*** |
Pairwise comparisons | |||
1 - 2 | -0.077 | 0.315 | 0.807 |
1 - 3 | 0.361 | 0.332 | 0.278 |
1 - 4 | -0.389 | 0.319 | 0.223 |
2 - 3 | 0.437 | 0.314 | 0.163 |
2 - 4 | -0.313 | 0.300 | 0.298 |
3 - 4 | -0.750 | 0.316 | 0.018* |
Other fixed effects | |||
0.141 | 0.044 | 0.001*** | |
0.004 | 0.002 | 0.004** | |
-0.001 | 0.000 | 0.011* | |
0.585 | 0.233 | 0.012* | |
-0.130 | 0.050 | 0.009** | |
Random effect | 0.458 | 0.228 |
Table of pairwise comparisons for the linear mixed model for daily Survey 3 trust. The group means for the four conditions are 3.16, 3.083, 3.52, and 2.77. Significant differences exist in Conditions 3 and 4 (difference -0.75). Other significant fixed effects are (0.141), (0.004), the interaction of and (-0.001), (0.585), and the interaction of and (-0.13).
For the daily Survey 3 reliance indicator, the best-fitting LMM including the condition had as fixed effects, and thus only depended on (1) the condition, (2) the perceived accuracy of the tool’s estimate, and (3) the perceived accuracy of the participant’s expectations. We did not include the trust score in this model, as it harmed the AIC and we found previously that it had no significant effect. Note the lack of interactions involving the condition. Group means, pairwise comparisons, and other coefficients for both the LMM and GLMM are shown in Table 6.
LMM | GLMM | ||||
Coefficient | |||||
Group means | |||||
Condition 1 | 0.324 | 0.111 | 0.003** | -0.476 | 0.421 |
Condition 2 | 0.595 | 0.140 | 0.000*** | 0.691 | 0.557 |
Condition 3 | 0.503 | 0.147 | 0.001*** | 0.333 | 0.574 |
Condition 4 | 0.594 | 0.141 | 0.000*** | 0.442 | 0.533 |
Pairwise comparisons | |||||
1 - 2 | 0.272 | 0.130 | 0.037* | 1.167 | 0.484 |
1 - 3 | 0.179 | 0.137 | 0.191 | 0.809 | 0.497 |
1 - 4 | 0.271 | 0.132 | 0.041* | 0.917 | 0.468 |
1 - 3 | -0.092 | 0.129 | 0.473 | -0.358 | 0.529 |
2 - 4 | -0.001 | 0.123 | 0.992 | -0.249 | 0.507 |
3 - 4 | 0.091 | 0.131 | 0.486 | 0.108 | 0.521 |
Other fixed effects | |||||
0.217 | 0.073 | 0.003** | |||
0.070 | 0.066 | 0.293 | |||
Random effect | 0.068 | 0.079 | -43.000 | 0.013 |
Table of pairwise comparisons for the linear mixed model and generalised linear mixed model for daily Survey 3 reliance. The group means for the four conditions are 0.324, 0.595, 0.503, and 0.594 for the LMM, and -0.476, 0.691, 0.333, and 0.442 for the GLMM. For the LMM, significant differences exist in Conditions 1 and 2 (difference 0.272) and Conditions 1 and 4 (difference 0.271). The fixed effect for (0.217) is also significant.
For the post-interaction Survey 4 trust score, the best-fitting OLS model including the condition was , and thus only depended on (1) the condition and (2) the perceived accuracy of the tool’s estimate on the first day. While we found no significant differences in trust scores between conditions, Condition 1 had higher trust scores than Condition 2, 3, and 4. Group means, pairwise comparisons, and other coefficients are shown in Table 7.
Coefficient | |||
Group means | |||
Condition 1 | 3.586 | 0.384 | 0.000*** |
Condition 2 | 3.273 | 0.445 | 0.000*** |
Condition 3 | 3.519 | 0.470 | 0.000*** |
Condition 4 | 3.211 | 0.432 | 0.000*** |
Pairwise comparisons | |||
1 - 2 | -0.313 | 0.345 | 0.364 |
1 - 3 | -0.067 | 0.385 | 0.862 |
1 - 4 | -0.374 | 0.338 | 0.268 |
2 - 3 | 0.246 | 0.345 | 0.475 |
2 - 4 | -0.061 | 0.292 | 0.834 |
3 - 4 | -0.308 | 0.338 | 0.363 |
Other coefficients | |||
0.4972 | 0.325 | 0.137 |
Table of pairwise comparisons for the ordinary least squares model for post-interaction Survey 4 trust. The group means for the four conditions are 3.586, 3.273, 3.519, and 3.211. Significant differences do not exist.
Our hypotheses did not capture the reality that different framings of uncertainty do not uniformly impact trust and reliance. Based on the daily Survey 3 measurements, Conditions 2 and 4 had the lowest reported trust ( for each condition: ) yet significantly higher reliance than Condition 1 ( for pairwise comparisons; for each condition: ). This observation can likely be attributed to the fact that we observed much less attrition in these conditions. It is possible that this was the result of survivorship bias, where participants in Conditions 1 and 3 chose to drop out instead of reporting lower trust scores. Yet, based on responses for Day 1 (when dropout rates were still relatively uniform), Condition 1 and 3 began with higher trust scores ( for each condition: ). Thus, adding either granularity or hedging to estimates may have helped participants to maintain interest in relying on the tool for their decision-making, even if their baseline trust was lower. Such designs may have augmented reliance by calibrating perceptions of its trustworthiness to be realistic but not necessarily optimal (Ferrario and Loi, 2022; Schemmer et al., 2023).
With respect to the effect of ranged estimates, Condition 3 resulted in a significantly higher pre-interaction Survey 2 trust score than Condition 1 (, for Conditions 1/3). This improvement is contrary to the prediction of Hypothesis 4, and was likely due to participants perceiving the ranges as adding useful information at least initially (Section 6.2.2). Compared to the findings of (Prabhudesai et al., 2023), who suggested that presenting uncertainty mitigates undue reliance, we hypothesise that the more task-aligned presentation of uncertainty in our context (direct estimates that abstract away probabilities) led to participants trusting the tool more. While Condition 3’s daily Survey 3 trust score was also higher than Condition 1, the difference was not significant (; for Conditions 1/3), and Condition 3’s post-interaction Survey 4 trust score was lower (but not significantly) than Condition 1 (; for Conditions 1/3). We also note that Condition 3’s trust scores were significantly higher than Condition 4 for both Survey 2 (; for Conditions 3/4) and Survey 3 (; for Conditions 3/4). Hedging may have increased error visibility to the extent that it offset the utility of the ranges, leading participants to think more critically about the estimated earnings and recommendations (as found by (Prabhudesai et al., 2023)) and thus decreasing trust according to Hypothesis 4.
Confounding Variables
Beyond the analyses described above, we performed additional analyses to assess the effect of possible confounding variables on our measurements. To do so, we added variables to the best-fitting models for these measurements that represented the constraint variant (maximum hours or minimum earnings/daily or weekly), the participant’s city (Los Angeles, New York, Chicago, or Houston), and the participant’s level of motivation (as measured by Question 5 in the formative Survey 1, “When I drive, it’s important to me that I make some minimum amount of money”; see Section A.1). We mostly did not find significant differences in trust and reliance measures between the different levels of these variables. There were two exceptions. First, participants who set maximum hours per day had significantly higher reliance for the daily Survey 3 (). We believe that this result can be partially attributed to the smaller number of participants for this variant, as well as one participant who consistently indicated their desire to increase their reliance on the tool every day. Second, participants who were more motivated to make a minimum amount of money had significantly higher trust for the pre-interaction Survey 2 (), but we did not observe any longitudinal effects for this variable. Our full quantitative results for these analyses are detailed in Appendix C.
6.2.2. Qualitative Results
In the following, we denote our interview participants as F1–F7 following the ordering in Section 6.1.3, where we also list full demographic details. Again, F1, F2, and F6 were placed in Condition 2; F5 was placed in Condition 3; and F3, F4, and F7 were placed in Condition 4.
Motivations and Routines
Participants reported a diversity of motivations and routines. F1 and F2 viewed gig driving as a primary source of income, while F3–F6 used it to supplement other sources of income. F1, F4, and F6 maintained consistent driving schedules; F3 and F5 scheduled their driving around other commitments; and F2 and F6 adjusted their schedules based on the level of demand and their own motivations. Meanwhile, F7 differed from other participants in that he only performed deliveries only during his free time and using a bicycle instead of driving. He felt that the tool did not capture the effects of these habits on his potential earnings.
Tool Benefits and Concerns
Except for F4 and F7, all participants felt that the time they spent interacting with the tool was worthwhile. F1–F3, F5, and F6 felt that the tool helped to visualise information about potential earnings in a helpful manner. This addresses a key pain point of gig drivers: F1–F6 had all reported challenges in estimating their potential earnings as a consequence of unpredictability in gig demand, pay, or location, or of gig platforms providing insufficient information. F2, F4, and F6 had also indicated that their earnings were impacted by the variability of bonuses and tips. In terms of other benefits, F1, F2, F5, and F6 felt that the tool provided more or different information, either relative to gig platforms or to Gridwise, and F5 and F6 found that the tool helped them to decide when to work. F4 was unconvinced of the utility of the tool due to the many factors that could influence earnings.
Constraint Design
In terms of alignment, F1, F3, F5, and F6 all felt that the constraints the tool allowed them to input mostly aligned with their goals and routines. F5 and F6 particularly appreciated the hourly granularity of the availability constraints. However, F1 and F3 surfaced a previously unknown design issue: it was not intuitive to them that they could select their availability individually for each time slot, and thus the constraints they had entered were suboptimal. Participants also exhibited perceptions of control over the tool; F1–F3, F5, and F6 understood how the constraints were used to generate the recommended schedules, while F3–F6 felt that the recommended schedules aligned with their constraints. F3 and F4, however, did not initially notice that the tool was selecting a subset of their available hours in its recommendations, likely due to the more restricted availability that they had entered.
Reliance and Compliance
The extent to which participants relied on and complied with the tool was also variable. In particular, we did not observe any consistency in patterns of behaviour even within the same condition. For reliance, F2 and F3 relied on the tool’s recommendations to the extent that they actively scheduled their day around reviewing the daily schedules; F5 and F6 viewed the tool more as a reference that they sometimes incorporated into their planning; and F1, F4, and F7 did not rely on the tool’s recommendations much. For compliance, F1–F3, F5, and F6 reported having adjusted their schedules in response to the recommendations at some points (with F5 in particular striving to perfectly follow the recommendations on weekdays but not on weekends), while F1, F2, F4, and F6 also reported instances where they maintained their existing schedules over the tool’s recommendations.
F1–F6 all reported that the extent to which they were motivated to comply with the recommended schedule depended upon how the tool’s estimates compared to their expectations or actual earnings. For instance, F3 decided to stop following the schedule on what she felt was a “slow day”; F4 was intrigued by optimistic hourly estimates that were much higher than his personal goal for those times, despite mostly choosing not to follow the schedule; F5 became more motivated to follow the schedules after observing an increase in earnings of 20–30%; and F6 chose not to follow the schedule for hours where he felt that the tool was underestimating potential earnings. In terms of longitudinal effects, F1, F2, and F5 indicated that the outcomes of their first one or two days of interaction impacted their willingness to follow the recommendations for the rest of the study, again in line with the findings of (Mou and Cohen, 2017; Mou et al., 2017). Meanwhile, F4 stated that he lost trust in the system after seeing that its estimates did not align with his expectations.
Perceptions of Uncertainty
One of the causes for participants selectively complying with the recommended schedules at some times but not others was the variability in the accuracy of its estimates. F1–F6 mentioned cases where the hourly earning estimates fell below their expected or actual earnings; F1–F4 and F6 mentioned cases where they fell above; and F5 and F6 mentioned cases where they were similar. However, this variability was less evident in daily and weekly estimates: in Condition 2, F1, F2, and F6 felt that their daily earnings were accurately estimated, and F1 and F2 also felt that their weekly earnings were accurately estimated. Regardless, F1, F3, and F5 indicated that they used the tool’s estimates as goals for their earnings. Some correctly recognised that external factors could impact the tool’s accuracy: for F1 and F3, their own behaviour; and for F3 and F6, inherent environmental uncertainty. F2 and F6 believed that the tool estimated their base earnings before bonuses and tips, whereas it actually estimated their total earnings.
Among participants in the conditions with ranges (Conditions 3 and 4), F3–F5 all liked its presence, although F7 did not. F3–F5 also all reported focusing on the optimistic estimate over the mean estimate to some extent. However, a major cause of F4 losing trust was that he considered the pessimistic estimates to be too low; F5 echoed this sentiment but otherwise had a positive impression. When prompted regarding whether they thought a driver other than themselves (e.g., F7) could have earned similarly to the pessimistic estimate, both F4 and F5 found it difficult to fathom, despite the hedging text presented to F4 by Condition 4. Among participants in the conditions without ranges (Condition 2), F1 and F2 indicated that they would have appreciated the range, while F6 suggested that the range would have had limited value since he was already viewing the mean estimates as being potentially pessimistic.
Suggestions for Improvement
Participants suggested a number of features that could improve the tool’s utility and trustworthiness. These included the ability to understand how bonuses, tips (F2, F6), and events (F5) affect earnings; to estimate driver supply (F2) and trip mileage (F3); to restrict or view estimates for different regions of their cities (F1, F6); to incorporate pre-scheduled gigs into the recommended schedule (F6); and to be able to integrate the tool more smoothly with gig platforms, other assistants such as Gridwise, or personal calendar apps (F1, F5, F6).
7. Limitations
Our central motivation in this work was to assess the effects of different framings of uncertainty on trust and reliance in an AI decision aid. To this end, we attempted to isolate the effect of this design dimension from others by first iterating on our tool’s design. In particular, we used two separate pilots to identify and implement a number of design desiderata based on participant feedback. However, we certainly cannot claim that the design of our tool was optimal for engendering trust, as also reflected in feedback from our interview participants (Section 6.2.2). Data availability limitations, for instance, meant that our tool was unable to provide more direct feedback on how participants’ performance compared to the tool’s estimates on a daily basis. We cannot exclude the possibility that some of these design choices had compounding impacts on our measurements of trust and reliance.
Also, despite our best efforts, we were only able to recruit a small sample of drivers from the four cities and six platforms for which we could access historical data. These individuals were at least aware, if not active users, of the Gridwise app, and thus they may have been more focused on the outcomes of their driving compared to the general gig driver population. One benefit of recruiting from this population is that we received high-quality responses and experienced relatively low attrition, which facilitated our analysis. However, we envision that a future study aimed at a broader population could uncover an additional wealth of insights, especially considering the variation that existed even in our small sample.
8. Conclusion
In this work, we investigated how different framings of uncertain estimates impacted the extent to which users were willing to trust in and rely on an AI decision aid within a context characterised by uncertainty and repeated interactions. As a case study, we deployed these framings as part of a longitudinal trial of a schedule recommendation tool to 51 gig drivers. Our quantitative results affirm the influence of users’ perceptions of the AI’s accuracy in estimating outcomes on trust and reliance, and demonstrate that framings of uncertainty have varying impacts on trust and reliance. While increased granularity and hedging language did not appear to improve trust like ranged estimates did, they did lead to participants expressing higher reliance and exhibiting lower attrition.
The most prominent conclusion we drew from the interviews we conducted, however, is that there is no one-size-fits-all approach to building trust. Even within the same condition, participants exhibited great variability in how they incorporated the AI system’s recommendations into their planning, and in how they reacted to the outcomes of their reliance. We thus echo calls to adopt user-centred design approaches for deployments of AI systems (Rantavuo, 2019; Wang et al., 2019). In particular, it is not enough to consider users as a homogeneous group. Designing trustworthy AI decision aids requires recognising the diverse needs and perceptions of different groups of users, and aligning explanations to the expectations of each group. Ideally, future empirical work should control more precisely for different user factors to support the identification of these groups. However, we also recognise that this is likely to be difficult: longitudinal effects such as those we uncovered imply that users’ latent attitudes cannot be easily assessed in a prospective manner, and users’ stated and revealed attitudes may also differ (Fietta et al., 2021).
Among the characteristics of users that are likely to shape their trust in a specific AI decision aid is their innate propensity to trust AI systems in general (Solberg et al., 2022). Various empirical studies have found users to exhibit either propensities towards algorithmic aversion (Mahmud et al., 2022) — a reluctance to rely on algorithms — or its polar opposite, algorithmic appreciation (Logg et al., 2019). These results are less likely to be contradictory than reflective of genuine attitudes that exist in different segments of the population; (Stein et al., 2024), for instance, found that algorithmic appreciation correlated with personality traits and age. Our longitudinal results support findings in single-shot settings that the context-specific expression of these general attitudes is influenced by framing (Hou and Jung, 2021) and outcomes (Dietvorst et al., 2015). But to what extent do specific interactions with AI systems conversely influence users’ general attitudes? We believe that this question merits further exploration.
More fundamentally, what should an AI decision aid’s goal be in the presence of uncertainty? Should it be to ensure the user’s compliance or simply their reliance? We demonstrate that reliance and compliance are not always aligned, and therefore that a trade-off must be made in certain contexts. However, we concur with (Schemmer et al., 2023) in that, if the nominal purpose of an AI decision aid is to support the user’s own goals, then a notion of appropriate reliance should take precedence over strict compliance due to the potential for misalignment in objectives. Since both the behaviour (Carroll et al., 2022) and presentation (Vasconcelos et al., 2023) of AI systems can both impact the reliance-compliance trade-off in users, we believe that careful requirements engineering is necessary to design AI decision aids with practical utility.
We viewed gig driving as a representative context that could provide rich data for our study. However, the challenges of uncertainty and repeated interactions are also prevalent in other deployment contexts of AI decision aids, such as healthcare, education, and criminal justice. We believe that our findings are sufficiently general that they could readily apply to other contexts; yet, the idiosyncrasies of different contexts may also influence the exact mechanisms by which framings of uncertainty impact trust and reliance. For instance, unlike how gig drivers receive near-instant feedback through earnings in our context, medical practitioners may not observe the impacts of their decisions until long afterwards. Our study is far from being the final word on how the framing of uncertainty impacts trust, but it suggests a fruitful line of future work tailored towards different deployment contexts of AI systems.
Acknowledgements.
We thank Joshua Mateer, Jenny T. Liang, Jose M. del Alamo, Ian Yang, Chenyang Yang, Stephanie Milani, and Niklas Kühl for their invaluable feedback and discussion on the study’s instruments, protocol, analysis, and manuscript. Also, we thank Linda Moreci for assistance with participant compensation, and Akshara Khare and Anagha Ravi Shankara for their contributions to the recommendation tool’s implementation. This work was conducted in collaboration with Gridwise; we thank Kunal Prajapati, Karan Moudgil, Ryan Green, and Brian Finnamore for their assistance with recruitment and data procurement. We were also supported by a research grant from Mobility21 (a US DOT National University Transportation Center) and the Tang Family Endowed Innovation Fund.References
- (1)
- Aoki (2021) Naomi Aoki. 2021. The importance of the assurance that “humans are still in the decision loop” for public trust in artificial intelligence: Evidence from an online experiment. Computers in Human Behavior 114 (2021), 106572.
- Bansal et al. (2019) Gagan Bansal, Besmira Nushi, Ece Kamar, Walter S. Lasecki, Daniel S. Weld, and Eric Horvitz. 2019. Beyond Accuracy: The Role of Mental Models in Human-AI Team Performance. In Proceedings of the 7th AAAI Conference on Human Computation and Crowdsourcing (HCOMP ’19). AAAI, Stevenson, USA, 2–11.
- Bansal et al. (2021) Gagan Bansal, Tongshuang Wu, Joyce Zhou, Raymond Fok, Besmira Nushi, Ece Kamar, Marco Tulio Ribeiro, and Daniel Weld. 2021. Does the Whole Exceed its Parts? The Effect of AI Explanations on Complementary Team Performance. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (CHI ’21). ACM, Yokohama, Japan, 1–16.
- Battifarano and Qian (2019) Matthew Battifarano and Sean Qian. 2019. Predicting real-time surge pricing of ride-sourcing companies. Transportation Research Part C: Emerging Technologies 107 (2019), 444–462.
- Beede et al. (2020) Emma Beede, Elizabeth Baylor, Fred Hersch, Anna Iurchenko, Lauren Wilcox, Paisan Ruamviboonsuk, and Laura M. Vardoulakis. 2020. A Human-Centered Evaluation of a Deep Learning System Deployed in Clinics for the Detection of Diabetic Retinopathy. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (CHI ’20). ACM, Honolulu, USA, 1–12.
- Berger et al. (2020) Thor Berger, Carl Benedikt Frey, Guy Levin, and Santosh Rao Danda. 2020. Uber happy? Work and well-being in the ‘Gig Economy’. Economic Policy 34, 99 (2020), 429–477.
- Bhatt et al. (2021) Umang Bhatt, Javier Antorán, Yunfeng Zhang, Q. Vera Liao, Prasanna Sattigeri, Riccardo Fogliato, Gabrielle Melançon, Ranganath Krishnan, Jason Stanley, Omesh Tickoo, Lama Nachman, Rumi Chunara, Madhulika Srikumar, Adrian Weller, and Alice Xiang. 2021. Uncertainty as a Form of Transparency: Measuring, Communicating, and Using Uncertainty. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (AIES ’21). ACM, Virtual, 401–413.
- Burgess et al. (2023) Eleanor R. Burgess, Ivana Jankovic, Melissa Austin, Nancy Cai, Adela Kapuścińska, Suzanne T. Currie, J. Marc Overhage, Erika S. Poole, and Jofsh Kaye. 2023. Healthcare AI Treatment Decision Support: Design Principles to Enhance Clinician Adoption and Trust. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). ACM, Hamburg, Germany, 1–19.
- Cai et al. (2019a) Carrie J. Cai, Samantha Winter, David Steiner, Lauren Wilcox, and Michael Terry. 2019a. “Hello AI”: Uncovering the Onboarding Needs of Medical Practitioners for Human–AI Collaborative Decision-Making. Proceedings of the ACM on Human-Computer Interaction 3 (2019), 1–24.
- Cai et al. (2019b) Carrie J. Cai, Samantha Winter, David Steiner, Lauren Wilcox, and Michael Terry. 2019b. Reliability and Inter-rater Reliability in Qualitative Research: Norms and Guidelines for CSCW and HCI Practice. Proceedings of the ACM on Human-Computer Interaction 3 (2019), 1–23.
- Cameron et al. (2022) Chris Cameron, Jason Hartford, Taylor Lundy, and Kevin Leyton-Brown. 2022. The Perils of Learning Before Optimizing. In Proceedings of the 36th AAAI Conference on Artificial Intelligence (AAAI ’21). AAAI, Virtual, 3708–3715.
- Carroll et al. (2022) Micah D Carroll, Anca Dragan, Stuart Russell, and Dylan Hadfield-Menell. 2022. Estimating and Penalizing Induced Preference Shifts in Recommender Systems. In Proceedings of the 39th International Conference on Machine Learning (ICML ’22). PMLR, Hawaii, 2686–2709.
- Cau et al. (2023) Federico Maria Cau, Hanna Hauptmann, Lucio Davide Spano, and Nava Tintarev. 2023. Effects of AI and Logic-Style Explanations on Users’ Decisions under Different Levels of Uncertainty. ACM Transactions on Interactive Intelligent Systems Just Accepted (2023), 1–41.
- Cavanaugh and Neath (2019) Joseph E. Cavanaugh and Andrew A. Neath. 2019. The Akaike information criterion: Background, derivation, properties, application, interpretation, and refinements. Wiley Interdisciplinary Reviews: Computational Statistics 11, 3 (2019), e1460.
- Chen et al. (2023) Yiqun T. Chen, Katharina Reinecke, Angela D. R. Smith, and Alexandra To. 2023. Why, when, and from whom: considerations for collecting and reporting race and ethnicity data in HCI. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). ACM, Hamburg, Germany, 1–15.
- Chyung et al. (2018) Seung Youn Chyung, Megan Kennedy, and Ingrid Campbell. 2018. Evidence-Based Survey Design: The Use of Ascending or Descending Order of Likert-Type Response Options. Performance Improvement 57, 9 (2018), 9–16.
- Chyung et al. (2017) Seung Youn Chyung, Katherine Roberts, Ieva Swanson, and Andrea Hankinson. 2017. Evidence-Based Survey Design: The Use of a Midpoint on the Likert Scale. Performance Improvement 56, 10 (2017), 15–23.
- Curtain (2023) Colin Curtain. 2023. QualCoder. University of Tasmania. https://github.com/ccbogel/QualCoder/releases/tag/3.3
- Danner et al. (2020) Simon Danner, Matthias Pfromm, and Klaus Bengler. 2020. Does Information on Automated Driving Functions and the Way of Presenting It before Activation Influence Users’ Behavior and Perception of the System? Information 11, 1 (2020), 54.
- Dell et al. (2012) Nicola Dell, Vidya Vaidyanathan, Indrani Medhi, Edward Cutrell, and William Thies. 2012. “Yours is better!”: participant response bias in HCI. In Proceedings of the 2012 CHI Conference on Human Factors in Computing Systems (CHI ’12). ACM, Austin, USA, 1321–1330.
- Dietvorst et al. (2015) Berkeley J. Dietvorst, Joseph P. Simmons, and Cade Massey. 2015. Algorithm aversion: People erroneously avoid algorithms after seeing them err. Journal of Experimental Psychology: General 144, 1 (2015), 114–126.
- Django (2023) Django. 2023. Django 4.1. Django Software Foundation. https://docs.djangoproject.com/en/4.1
- Feng et al. (2023) K. J. Kevin Feng, Tony W. Li, and Amy X. Zhang. 2023. Understanding Collaborative Practices and Tools of Professional UX Practitioners in Software Organizations. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). ACM, Hamburg, Germany, 1–20.
- Ferrario and Loi (2022) Andrea Ferrario and Michele Loi. 2022. How Explainability Contributes to Trust in AI. In Proceedings of the 2022 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’22). ACM, Seoul, South Korea, 1457–1466.
- Ferraro et al. (2020) Andres Ferraro, Dietmar Jannach, and Xavier Serra. 2020. Exploring Longitudinal Effects of Session-based Recommendations. In Proceedings of the 14th ACM Conference on Recommender Systems (RecSys ’20). ACM, Virtual, 474–479.
- Fietta et al. (2021) Valentina Fietta, Francesca Zecchinato, Brigida Di Stasi, Mirko Polato, and Merylin Monaro. 2021. Dissociation Between Users’ Explicit and Implicit Attitudes Toward Artificial Intelligence: An Experimental Study. IEEE Transactions on Human-Machine Systems 52, 3 (2021), 481–489.
- Figma, Inc. (nd) Figma, Inc. n.d.. Figma. Figma, Inc. https://figma.com
- Gillath et al. (2021) Omri Gillath, Ting Ai, Michael S. Branicky, Shawn Keshmiri, Robert B. Davison, and Ryan Spaulding. 2021. Attachment and trust in artificial intelligence. Computers in Human Behavior 115 (2021), 106607.
- Gretzel and Fesenmaier (2006) Ulrike Gretzel and Daniel R. Fesenmaier. 2006. Persuasion in Recommender Systems. International Journal of Electronic Commerce 11, 2 (2006), 81–100.
- Griesbach et al. (2019) Kathleen Griesbach, Adam Reich, Luke Elliott-Negri, and Ruth Milkman. 2019. Algorithmic Control in Platform Food Delivery Work. Socius 5 (2019), 1–15.
- Ho et al. (2013) Geoffrey Ho, Nada Pavlovic, Vincent Myers, and Robert Arrabito. 2013. Reducing False Alarms in Automated Target Recognition by Lowering the Level of Automation. In Proceedings of the 2013 Human Factors and Ergonomics Society Annual Meeting (HFES ’13). SAGE, San Diego, USA, 374–378.
- Holmström and Hällgren (2022) Jonny Holmström and Markus Hällgren. 2022. AI management beyond the hype: exploring the co-constitution of AI and organizational context. AI & Society 37 (2022), 1575–1585.
- Hou and Jung (2021) Yoyo Tsung-Yu Hou and Malte F. Jung. 2021. Who is the Expert? Reconciling Algorithm Aversion and Algorithm Appreciation in AI-Supported Decision Making. Proceedings of the ACM on Human-Computer Interaction 5 (2021), 1–25.
- Jacobs et al. (2021) Maia Jacobs, Jeffrey He, Melanie F. Pradier, Barbara Lam, Andrew C. Ahn, Thomas H. McCoy, Roy H. Perlis, Finale Doshi-Velez, and Krzysztof Z. Gajos. 2021. Designing AI for Trust and Collaboration in Time-Constrained Medical Decisions: A Sociotechnical Lens. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (CHI ’21). ACM, Yokohama, Japan, 1–14.
- Jian et al. (2000) Jiun-Yin Jian, Ann M. Bisantz, and Colin G. Drury. 2000. Foundations for an Empirically Determined Scale of Trust in Automated Systems. International Journal of Cognitive Ergonomics 4, 1 (2000), 53–71.
- Jiang et al. (2004) Xiaochun Jiang, Mohammad T. Khasawneh, Reena Master, Shannon R. Bowling, Anand K. Gramopadhye, Brian J. Melloy, and Larry Grimes. 2004. Measurement of human trust in a hybrid inspection system based on signal detection theory measures. International Journal of Industrial Ergonomics 34, 5 (2004), 407–419.
- Kaplan et al. (2023) Alexandra D. Kaplan, Theresa T. Kessler, J. Christopher Brill, and P. A. Hancock. 2023. Trust in Artificial Intelligence: Meta-Analytic Findings. Human Factors 65, 2 (2023), 337–359.
- Kaufman and Kirsh (2022) Robert A. Kaufman and David J. Kirsh. 2022. Cognitive Differences in Human and AI Explanation. In Proceedings of the 44th Annual Conference of the Cognitive Science Society (CogSci ’22). eScholarship, Toronto, Canada, 2694–2700.
- Khan et al. (2022) Hassan Ali Khan, Muhammad Shahzad, Hassan Iqbal, and Guoliang Jin. 2022. RMS: Removing Barriers to Analyze the Availability and Surge Pricing of Ridesharing Services. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (CHI ’22). ACM, New Orleans, USA, 1–18.
- Ko et al. (2015) Amy J. Ko, Thomas D. LaToza, and Margaret M. Burnett. 2015. A practical guide to controlled experiments of software engineering tools with human participants. Empirical Software Engineering 20 (2015), 110–141.
- Kocielnik et al. (2019) Rafal Kocielnik, Saleema Amershi, and Paul N. Bennett. 2019. Will You Accept an Imperfect AI? Exploring Designs for Adjusting End-user Expectations of AI Systems. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems (CHI ’19). ACM, Glasgow, UK, 1–14.
- Kohn et al. (2021) Spencer C. Kohn, Ewart J. de Visser, Eva Wiese, Yi-Ching Lee, and Tyler H. Shaw. 2021. Measurement of Trust in Automation: A Narrative Review and Reference Guide. Frontiers in Psychology 12 (2021), 604977.
- Komiak and Benbasat (2006) Sherrie Y. X. Komiak and Izak Benbasat. 2006. The Effects of Personalization and Familiarity on Trust and Adoption of Recommendation Agents. MIS Quarterly 30, 4 (2006), 941–960.
- Kunkel et al. (2019) Johannes Kunkel, Tim Donkers, Lisa Michael, Catalin-Mihai Barbu, and Jürgen Ziegler. 2019. Let Me Explain: Impact of Personal and Impersonal Explanations on Trust in Recommender Systems. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems (CHI ’19). ACM, Scotland, UK, 1–12.
- Lee and Moray (1992) John Lee and Neville Moray. 1992. Trust, control strategies and allocation of function in human-machine systems. Ergonomics 10 (1992), 1243–1270.
- Lee and See (2004) John D. Lee and Katrina A. See. 2004. Trust in Automation: Designing for Appropriate Reliance. Human Factors 46, 1 (2004), 50–80.
- Lee et al. (2015) Min Kyung Lee, Danyel Kusbit, Evan Metsky, and Laura Dabbish. 2015. Working with Machines: The Impact of Algorithmic and Data-Driven Management on Human Workers. In Proceedings of the 2015 CHI Conference on Human Factors in Computing Systems (CHI ’15). ACM, Seoul, South Korea, 1–13.
- Liang and Willemsen (2022) Yu Liang and Martijn C. Willemsen. 2022. Exploring the longitudinal effects of nudging on users’ music genre exploration behavior and listening preferences. In Proceedings of the 16th ACM Conference on Recommender Systems (RecSys ’22). ACM, Seattle, USA, 3–13.
- Liao and Sundar (2022a) Mengqi Liao and S. Shyam Sundar. 2022a. When E-Commerce Personalization Systems Show and Tell: Investigating the Relative Persuasive Appeal of Content-Based versus Collaborative Filtering. Journal of Advertising 51 (2022), 256–267.
- Liao and Sundar (2022b) Q. Vera Liao and S. Shyam Sundar. 2022b. Designing for Responsible Trust in AI Systems: A Communication Perspective. In Proceedings of the 2022 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’22). ACM, Seoul, Korea, 1257–1268.
- Lockey et al. (2021) Steven Lockey, Nicole Gillespie, Daniel Holm, and Ida Asadi Someh. 2021. A Review of Trust in Artificial Intelligence: Challenges, Vulnerabilities and Future Directions. In Proceedings of the 54th Hawaii International Conference on System Sciences (HICSS ’21). AIS, Honolulu, USA, 5463–5472.
- Logg et al. (2019) Jennifer M. Logg, Julia A. Minson, and Don A. Moore. 2019. Algorithm appreciation: People prefer algorithmic to human judgment. Organizational Behavior and Human Decision Processes 151 (2019), 90–103.
- Ma and Hanrahan (2019) Ning F. Ma and Benjamin V. Hanrahan. 2019. Part-Time Ride-Sharing: Recognizing the Context in which Drivers Ride-Share and its Impact on Platform Use. Proceedings of the ACM on Human-Computer Interaction 3 (2019), 1–17.
- Madsen and Gregor (2000) Maria Madsen and Shirley Gregor. 2000. Measuring Human-Computer Trust. In Proceedings of the 11th Australasian Conference on Information Systems (ACIS ’00). AAIS, Brisbane, Australia, 6–8.
- Maffie (2023) Michael David Maffie. 2023. Visible hands: How gig companies shape workers’ exposure to market risk. Industrial Relations 63 (2023), 1–21.
- Mahmud et al. (2022) Hasan Mahmud, A.K.M. Najmul Islam, Syed Ishtiaque Ahmed, and Kari Smolander. 2022. What influences algorithmic decision-making? A systematic literature review on algorithm aversion. Technological Forecasting and Social Change 175 (2022), 121390.
- Mayer et al. (1995) Roger C. Mayer, James H. Davis, and F. David Schoorman. 1995. An Integrative Model of Organizational Trust. The Academy of Management Review 20, 3 (1995), 709–734.
- McConvey et al. (2023) Kelly McConvey, Shion Guha, and Anastasia Kuzminykh. 2023. A Human-Centered Review of Algorithms in Decision-Making in Higher Education. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). ACM, Hamburg, Germany, 1–15.
- McGregor (2021) Sean McGregor. 2021. Preventing Repeated Real World AI Failures by Cataloging Incidents: The AI Incident Database. In Proceedings of the 35th AAAI Conference on Artificial Intelligence (AAAI ’21). AAAI, Virtual, 15458–15463.
- Mcknight et al. (2011) D. Harrison Mcknight, Michelle Carter, Jason Bennett Thatcher, and Paul F. Clay. 2011. Trust in a specific technology: An investigation of its components and measures. ACM Transactions on Management Information Systems 2, 2 (2011), 1–25.
- Meske and Bunde (2020) Christian Meske and Enrico Bunde. 2020. Transparency and Trust in Human-AI-Interaction: The Role of Model-Agnostic Explanations in Computer Vision-Based Decision Support. In Proceedings of the 2020 International Conference on Human-Computer Interaction (HCII ’20). Springer, Copenhagen, Denmark, 54–69.
- Miao et al. (2023) Wei Miao, Yiting Deng, Wei Wang, Yongdong Liu, and Christopher S. Tang. 2023. The effects of surge pricing on driver behavior in the ride-sharing market: Evidence from a quasi-experiment. Journal of Operations Management 69, 5 (2023), 794–822.
- Milana and Ashta (2021) Carlo Milana and Arvind Ashta. 2021. Artificial intelligence techniques in finance and financial markets: A survey of the literature. Strategic Change 30, 3 (2021), 189–209.
- Mou and Cohen (2017) Jian Mou and Jason F. Cohen. 2017. Trust and online consumer health service success: A longitudinal study. Information Development 33, 2 (2017), 169–189.
- Mou et al. (2017) Jian Mou, Dong-Hee Shin, and Jason Cohen. 2017. Understanding trust and perceived usefulness in the consumer acceptance of an e-service: a longitudinal investigation. Behaviour & Information Technology 36, 2 (2017), 125–139.
- Naiseh et al. (2021) Mohammad Naiseh, Reem S. Al-Mansoori, Dena Al-Thani, Nan Jiang, and Raian Ali. 2021. Nudging through Friction: An Approach for Calibrating Trust in Explainable AI. In Proceedings of the 8th International Conference on Behavioral and Social Computing (BESC ’21). IEEE, Virtual, 1–5.
- Naiseh et al. (2023) Mohammad Naiseh, Dena Al-Thani, Nan Jiang, and Raian Ali. 2023. How the different explanation classes impact trust calibration: The case of clinical decision support systems. International Journal of Human-Computer Studies 169 (2023), 102941.
- Nourani et al. (2021) Mahsan Nourani, Chiradeep Roy, Jeremy E. Block, Donald R. Honeycutt, Tahrima Rahman, Eric D. Ragan, and Vibhav Gogate. 2021. Anchoring Bias Affects Mental Model Formation and User Reliance in Explainable AI Systems. In Proceedings of the 26th International Conference on Intelligent User Interfaces (IUI ’21). ACM, College Station, USA, 340–350.
- Prabhudesai et al. (2023) Snehal Prabhudesai, Leyao Yang, Sumit Asthana, Xun Huan, Q. Vera Liao, and Nikola Banovic. 2023. Understanding Uncertainty: How Lay Decision-makers Perceive and Interpret Uncertainty in Human-AI Decision Making. In Proceedings of the ACM Conference on Intelligent User Interfaces 2023 (IUI ’23). ACM, Sydney, Australia, 379–396.
- Rader et al. (2018) Emilee Rader, Kelley Cotter, and Janghee Cho. 2018. Explanations as Mechanisms for Supporting Algorithmic Transparency. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems (CHI ’18). ACM, Montréal, Canada, 1–14.
- Rantavuo (2019) Heli Rantavuo. 2019. Designing for intelligence: User-centred design in the age of algorithms. In Proceedings of the 5th International ACM In-Cooperation HCI and UX Conference (CHIuXiD ’23). ACM, Bali, Indonesia, 182–187.
- Riahi et al. (2021) Youssra Riahi, Tarik Saikouk, Angappa Gunasekaran, and Ismail Badraoui. 2021. Artificial intelligence applications in supply chain: A descriptive bibliometric analysis and future research directions. Expert Systems with Applications 173 (2021), 114702.
- Rosenblat and Stark (2016) Alex Rosenblat and Luke Stark. 2016. Algorithmic Labor and Information Asymmetries: A Case Study of Uber’s Drivers. International Journal Of Communication 10 (2016), 1–27.
- Russell et al. (2017) Stephen Russell, Ira S. Moskowitz, and Adrienne Raglin. 2017. Human Information Interaction, Artificial Intelligence, and Errors. In Autonomy and Artificial Intelligence: A Threat or Savior? Springer, Berlin, 71–101.
- Saldaña (2015a) Johnny Saldaña. 2015a. First-Cycle Coding Methods. In The Coding Manual for Qualitative Researchers (3rd ed.). SAGE, San Diego, USA, 67–206.
- Saldaña (2015b) Johnny Saldaña. 2015b. Second-Cycle Coding Methods. In The Coding Manual for Qualitative Researchers (3rd ed.). SAGE, San Diego, USA, 233–268.
- Saßmannshausen et al. (2021) Till Saßmannshausen, Peter Burggräf, Johannes Wagner, Marc Hassenzahl, Thomas Heupel, and Fabian Steinberg. 2021. Trust in artificial intelligence within production management — an exploration of antecedents. Ergonomics 64, 10 (2021), 1333–1350.
- Scharowski et al. (2023) Nicolas Scharowski, Michaela Benk, Swen J. Kühne, Léane Wettstein, and Florian Brühlmann. 2023. Certification Labels for Trustworthy AI: Insights From an Empirical Mixed-Method Study. In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’23). ACM, Chicago, USA, 248–262.
- Schemmer et al. (2023) Max Schemmer, Niklas Kühl, Carina Benz, Andrea Bartos, and Gerhard Satzger. 2023. Appropriate Reliance on AI Advice: Conceptualization and the Effect of Explanations. In Proceedings of the 28th International Conference on Intelligent User Interfaces (IUI ’23). ACM, Sydney, Australia, 410–422.
- Schoeffer et al. (2022) Jakob Schoeffer, Niklas Kuehl, and Yvette Machowski. 2022. “There Is Not Enough Information”: On the Effects of Explanations on Perceptions of Informational Fairness and Trustworthiness in Automated Decision-Making. In Proceedings of the 2022 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’22). ACM, Seoul, Korea, 1616–1628.
- Seabold and Perktold (2010) Skipper Seabold and Josef Perktold. 2010. statsmodels: Econometric and statistical modeling with Python. In Proceedings of the 9th Python in Science Conference (SciPy ’10). Python in Science Conference, Austin, USA, 92–96.
- Shin (2021) Donghee Shin. 2021. The effects of explainability and causability on perception, trust, and acceptance: Implications for explainable AI. International Journal of Human-Computer Studies 146 (2021), 102551.
- Singh et al. (2023) Anubha Singh, Patricia Garcia, and Silvia Lindtner. 2023. Old Logics, New Technologies: Producing a Managed Workforce on On-Demand Service Platforms. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). ACM, Hamburg, Germany, 1–15.
- Sivaraman et al. (2023) Venkatesh Sivaraman, Leigh A. Bukowski, Joel Levin, Jeremy M. Kahn, and Adam Perer. 2023. Ignore, Trust, or Negotiate: Understanding Clinician Acceptance of AI-Based Treatment Recommendations in Health Care. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). ACM, Hamburg, Germany, 1–18.
- Skjuve et al. (2023) Marita Skjuve, Asbjørn Følstad, and Petter Bae Brandtzaeg. 2023. The User Experience of ChatGPT: Findings from a Questionnaire Study of Early Users. In Proceedings of the 5th International Conference on Conversational User Interfaces (CUI ’23). ACM, New York, USA, 1–10.
- Solberg et al. (2022) Elizabeth Solberg, Magnhild Kaarstad, Maren H. Rø Eitrheim, Rossella Bisio, Kine Reegard, and Marten Bloch. 2022. A Conceptual Model of Trust, Perceived Risk, and Reliance on AI Decision Aids. Group & Organization Management 47, 2 (2022), 187–222.
- Stein et al. (2023) Jake M L Stein, Vidminas Vizgirda, Max Van Kleek, Reuben Binns, Jun Zhao, Rui Zhao, Naman Goel, George Chalhoub, Wael S Albayaydh, and Nigel Shadbolt. 2023. ‘You are you and the app. There’s nobody else.’: Building Worker-Designed Data Institutions within Platform Hegemony. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). ACM, Hamburg, Germany, 1–26.
- Stein et al. (2024) Jan-Philipp Stein, Tanja Messingschlager, Timo Gnambs, Fabian Hutmacher, and Markus Appel. 2024. Attitudes towards AI: measurement and associations with personality. Scientific Reports 14 (2024), 2909.
- Stuck et al. (2021) Rachel E. Stuck, Brittany E. Holthausen, and Bruce N. Walker. 2021. Chapter 8 - The role of risk in human-robot trust. In Trust in Human-Robot Interaction. Academic Press, Cambridge, 179–194.
- Sun et al. (2023) Yuan Sun, Magdalayna Drivas, Mengqi Liao, and S. Shyam Sundar. 2023. When Recommender Systems Snoop into Social Media, Users Trust them Less for Health Advice. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). ACM, Hamburg, Germany, 1–14.
- Sundar et al. (2008) S. Shyam Sundar, Anne Oeldorf-Hirsch, and Qian Xu. 2008. The Bandwagon Effect of Collaborative Filtering Technology. In Proceedings of the 2008 CHI Conference on Human Factors in Computing Systems (CHI ’08). ACM, Florence, Italy, 3453–3458.
- Suwanaposee et al. (2023) Pang Suwanaposee, Carl Gutwin, Zhe Chen, and Andy Cockburn. 2023. ‘Specially For You’ — Examining the Barnum Effect’s Influence on the Perceived Quality of System Recommendations. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). ACM, Hamburg, Germany, 1–11.
- Tan et al. (2021) Zhi Ming Tan, Nikita Aggarwal, Josh Cowls, Jessica Morley, Mariarosaria Taddeo, and Luciano Floridi. 2021. The ethical debate about the gig economy: A review and critical analysis. Technology and Society 65 (2021), 101594.
- Textor et al. (2022) Claire Textor, Rui Zhang, Jeremy Lopez, Beau G. Schelble, Nathan J. McNeese, Guo Freeman, Richard Pak, Chad Tossell, and Ewart J. de Visser. 2022. Exploring the Relationship Between Ethics and Trust in Human–Artificial Intelligence Teaming: A Mixed Methods Approach. Journal of Cognitive Engineering and Decision Making 16, 4 (2022), 252–281.
- Tomsett et al. (2020) Richard Tomsett, Alun Preece, Dave Braines, Federico Cerutti, Supriyo Chakraborty, Mani Srivastava, Gavin Pearson, and Lance Kaplan. 2020. Rapid Trust Calibration through Interpretable and Uncertainty-Aware AI. Patterns 1, 4 (2020), 100049.
- Ueno et al. (2022) Takane Ueno, Yuto Sawa, Yeongdae Kim, Jacqueline Urakami, Hiroki Oura, and Katie Seaborn. 2022. Trust in Human-AI Interaction: Scoping Out Models, Measures, and Methods. In Extended Abstracts of the 2022 CHI Conference on Human Factors in Computing Systems (CHI EA ’22). ACM, New Orleans, USA, 1–7.
- van Berkel and Kostakos (2021) Niels van Berkel and Vassilis Kostakos. 2021. Recommendations for Conducting Longitudinal Experience Sampling Studies. In Advances in Longitudinal HCI Research. Springer, Berlin, 59–78.
- Vasconcelos et al. (2023) Helena Vasconcelos, Matthew Jörke, Madeleine Grunde-McLaughlin, Tobias Gerstenberg, Michael S. Bernstein, and Ranjay Krishna. 2023. Explanations Can Reduce Overreliance on AI Systems During Decision-Making. Proceedings of the ACM on Human-Computer Interaction 7 (2023), 1–38.
- Wang et al. (2021) Dakuo Wang, Liuping Wang, Zhan Zhang, Ding Wang, Haiyi Zhu, Yvonne Gao, Xiangmin Fan, and Feng Tian. 2021. “Brilliant AI Doctor” in Rural Clinics: Challenges in AI-Powered Clinical Decision Support System Deployment. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (CHI ’21). ACM, Yokohama, Japan, 1–18.
- Wang et al. (2019) Danding Wang, Qian Yang, Ashraf Abdul, and Brian Y. Lim. 2019. Designing Theory-Driven User-Centric Explainable AI. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems (CHI ’19). ACM, Glasgow, UK, 1–15.
- Wang and Yin (2021) Xinru Wang and Ming Yin. 2021. Are Explanations Helpful? A Comparative Study of the Effects of Explanations in AI-Assisted Decision-Making. In Proceedings of the 26th International Conference on Intelligent User Interfaces (IUI ’21). ACM, College Station, USA, 318–328.
- Wintersberger et al. (2021) Philipp Wintersberger, Frederica Janotta, Jakob Peintner, Andreas Löcken, and Andreas Riener. 2021. Evaluating feedback requirements for trust calibration in automated vehicles. it — Information Technology 63, 2 (2021), 1–12.
- Wood et al. (2019) Alex J Wood, Mark Graham, Vili Lehdonvirta, and Isis Hjorth. 2019. Good Gig, Bad Gig: Autonomy and Algorithmic Control in the Global Gig Economy. Employment and Society 33, 1 (2019), 56–75.
- Wu et al. (2019) Qingjun Wu, Hao Zhang, Zhen Li, and Kai Liu. 2019. Labor control in the gig economy: Evidence from Uber in China. Journal of Industrial Relations 61, 4 (2019), 574–596.
- Yang et al. (2023) Qian Yang, Yuexing Hao, Kexin Quan, Stephen Yang, Yiran Zhao, Volodymyr Kuleshov, and Fei Wang. 2023. Harnessing Biomedical Literature to Calibrate Clinicians’ Trust in AI Decision Support Systems. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). ACM, Hamburg, Germany, 1–14.
- Završnik (2019) Aleš Završnik. 2019. Algorithmic justice: Algorithms and big data in criminal justice settings. European Journal of Criminology 18, 5 (2019), 623–642.
- Zhang et al. (2023) Angie Zhang, Alexander Boltz, Jonathan Lynn, Chun-Wei Wang, and Min Kyung Lee. 2023. Stakeholder-Centered AI Design: Co-Designing Worker Tools with Gig Workers through Data Probes. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). ACM, Hamburg, Germany, 1–18.
- Zhang et al. (2022) Angie Zhang, Alexander Boltz, Chun-Wei Wang, and Min Kyung Lee. 2022. Algorithmic Management Reimagined For Workers and By Workers: Centering Worker Well-Being in Gig Work. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (CHI ’22). ACM, New Orleans, USA, 1–20.
- Zhang et al. (2020) Yunfeng Zhang, Q. Vera Liao, and Rachel K. E. Bellamy. 2020. Effect of confidence and explanation on accuracy and trust calibration in AI-assisted decision making. In Proceedings of the 2020 Conference on Fairness, Accountability, and Transparency (FAccT ’20). ACM, Barcelona, Spain, 295–305.
- Zhou et al. (2021) Meizi Zhou, Jingjng Zhang, and Gediminas Adomavicius. 2021. Longitudinal Impact of Preference Biases on Recommender Systems’ Performance. Technical Report 2021-10. Kelley School of Business Research Paper.
Appendix A Survey Questions
A.1. Survey 1 — Formative Questions
1.
Please tell us which of the following regions you primarily drive in.
Los Angeles
New York
Chicago
Houston
2.
Please tell us which of the following services you currently drive for.
DoorDash
Grubhub
Instacart
Lyft
Uber
Uber Eats
Other
3.
In 2 or 3 sentences, please tell us what you like most about driving for ridesharing and/or delivery services.
4.
In 2 or 3 sentences, please tell us what you like least about driving for ridesharing and/or delivery services.
How much do you agree or disagree with the following statements?
5.
When I drive, it’s important to me that I make some minimum amount of money.
Strongly disagree
Disagree
Not sure
Agree
Strongly agree
6.
When I drive, I have an accurate sense of how much money I will make.
Strongly disagree
Disagree
Not sure
Agree
Strongly agree
7.
When I don’t earn the amount that I expect to from driving, it causes difficulties for me.
Strongly disagree
Disagree
Not sure
Agree
Strongly agree
8.
I try to stick to a regular routine for times and places to drive.
Strongly disagree
Disagree
Not sure
Agree
Strongly agree
9.
I am happy with how I currently decide when and where to drive.
Strongly disagree
Disagree
Not sure
Agree
Strongly agree
10.
Please tell us your age.
11.
Please tell us what gender you identify as.
Male
Female
Non-binary
Prefer not to disclose
Other
12.
Please tell us your highest education level.
Less than high school
High school
Some two-year professional degree
Some undergraduate degree
Some graduate degree (MS, PhD, JD, or MD)
A.2. Survey 2 — Pre-Interaction Trust
How much do you agree or disagree with the following statements?
1.
I understand how the tool used my answers to generate this recommended schedule.
Strongly disagree
Disagree
Not sure
Agree
Strongly agree
This item is based on item U2 from the Perceived Understandability questions of (Madsen and Gregor, 2000), “I understand how the system will assist me with decisions I have to make.” As the constraint page is intended to encapsulate the user’s decision-making process, we consider the generation of the recommended schedule to be how the tool assists the user with their decisions.
2.
I feel that I can rely on the tool to produce recommendations which accommodate the things that matter most to me.
Strongly disagree
Disagree
Not sure
Agree
Strongly agree
This item is based on item R4 from the Perceived Reliability questions of (Madsen and Gregor, 2000), “I can rely on the system to function properly.” We consider the tool to be properly functioning if its recommendations account for the user’s goals and preferences.
3.
I feel that the driving times recommended by the tool are as good as what an experienced driver would recommend to me.
Strongly disagree
Disagree
Not sure
Agree
Strongly agree
This item is based on item T3 from the Perceived Technical Competence questions of (Madsen and Gregor, 2000), “The advice the system produces is as good as that which a highly competent person could produce.” In our context, the tool’s advice is its recommended schedule of driving times, and to our participants a competent individual would be an experienced driver.
4.
I feel that the times suggested by the tool are good even if I don’t know for certain that they will maximise my earnings / minimise my hours [Depending on the constraints selected].
Strongly disagree
Disagree
Not sure
Agree
Strongly agree
This item is based on item F1 from the Faith questions of (Madsen and Gregor, 2000), “I believe advice from the system even when I don’t know for certain that it is correct.” Again, in our context, the tool’s advice is its recommended schedule of driving times. Since we cannot apply a clear notion of correctness to continuous estimates of earnings, we reworded this question to focus on alignment with the user’s objectives.
5.
I would like to use the tool to decide my driving hours in the future.
Strongly disagree
Disagree
Not sure
Agree
Strongly agree
This item is based on item P4 from the Personal Attachment questions of (Madsen and Gregor, 2000), “I like using the system for decision making.” We reworded the question to better assess participants’ level of intended future reliance on the tool.
A.3. Survey 3 — Daily Trust
1.
How often did you follow the times in the recommended schedule today?
I did not follow the recommendations at all
I followed the recommendations for one hour during the day
I followed the recommendations for two or three hours during the day
I followed the recommendations for four or more hours during the day
2.
How satisfied do you feel you are with your earnings from today?
Very dissatisfied
Somewhat dissatisfied
Neither satisfied nor dissatisfied
Somewhat satisfied
Very satisfied
3.
As far as you remember, how did your earnings today compare to your expectations?
Lower
About the same
Higher
Not sure
4.
As far as you remember, how did your earnings today compare to the tool’s estimate?
Lower
About the same
Higher
Not sure
How much do you agree or disagree with the following statements?
5.
I felt that the recommended schedule provided by the tool was easy to follow.
Strongly disagree
Disagree
Not sure
Agree
Strongly agree
This item is based on item U4 of the Perceived Understandability questions of (Madsen and Gregor, 2000), “It is easy to follow what the system does.” Instead of asking the user about the tool’s operation generally, we focused the question on the interpretability of its recommended schedule for that day.
6.
I felt that the recommended schedule provided all of the information that I needed to decide when to drive.
Strongly disagree
Disagree
Not sure
Agree
Strongly agree
This item is based on item R1 of the Perceived Reliability questions of (Madsen and Gregor, 2000), “The system always provides the advice I require to make my decision.” Again, in our context, the tool’s advice is its recommended schedule. We focused the question on the user’s decisions for that particular day.
7.
When I was unsure of when to drive today, I followed the recommended schedule.
Strongly disagree
Disagree
Not sure
Agree
Strongly agree
This item is based on item F2 from the Faith questions of (Madsen and Gregor, 2000), “When I am uncertain about a decision I believe the system rather than myself.” We focused the question on the user’s decisions for that particular day, and reworded “believe” to “follow” to assess compliance more clearly.
We left out questions based on Perceived Technical Competence and Personal Attachment for length.
8.
Which of the following statements do you agree with most?
I intend to rely on the tool less tomorrow than I did today
I intend to rely on the tool about the same tomorrow as I did today
I intend to rely on the tool more tomorrow than I did today
I intend to pause my interaction with the tool for one day tomorrow
A.4. Survey 4 — Post-Interaction Trust
How much do you agree or disagree with the following statements?
1.
I feel that I have become familiar with how to use the tool.
Strongly disagree
Disagree
Not sure
Agree
Strongly agree
This item is based on item 12 of (Jian et al., 2000)’s instrument, “I am familiar with the system.” We reworded the question in light of the fact that users did not have any existing experience with using the tool before the study.
2.
When I am using a navigation app that suggests routes to me, I feel like I would want to follow the suggestions more if the app asked me questions about my preferences (like this tool did) before giving its suggestions.
Strongly disagree
Disagree
Not sure
Agree
Strongly agree
This item is an original question that prompts the participant to consider their interactions with other types of recommendation systems. It assesses the extent to which participants would appreciate granular controls based on their preferences in such systems.
3.
Were there questions that you wanted the tool to ask you that it didn’t? If so, please tell us about them in 2 or 3 sentences.
How much do you agree or disagree with the following statements?
4.
I felt that I was able to trust the schedules recommended by the tool.
Strongly disagree
Disagree
Not sure
Agree
Strongly agree
This item is based on item 11 of (Jian et al., 2000)’s instrument, “I can trust the system.” We focused the scope of this question on the output of the tool, the recommended schedule, rather than the tool as a whole.
5.
I felt that I was able to depend on the schedules recommended by the tool for deciding when to drive.
Strongly disagree
Disagree
Not sure
Agree
Strongly agree
This item is based on item 9 of (Jian et al., 2000)’s instrument, “The system is dependable.” Again, we focused the scope of this question on the output of the tool, the recommended schedule.
6.
When I am using a navigation app that suggests routes to me, I feel like I would want to follow the suggestions more if the app gave me information about the minimum and maximum possible time of the trip (similar to what this tool did).
Strongly disagree
Disagree
Not sure
Agree
Strongly agree
This item is an original question that prompts the participant to consider their interactions with other types of recommendation systems. It assesses the extent to which participants would appreciate increased error visibility through ranged estimates in such systems.
7.
I felt that the recommended schedule was misleading.
Strongly disagree
Disagree
Not sure
Agree
Strongly agree
This item is based on item 1 of (Jian et al., 2000)’s instrument, “The system is deceptive.” In addition to focusing the scope of this question on the output of the tool, we also reworded “deceptive” to “misleading” to capture the broader possibility of the tool being perceived as unintentionally providing incorrect information.
8.
I felt that the recommended schedule harmed my earnings.
Strongly disagree
Disagree
Not sure
Agree
Strongly agree
This item is based on item 5 of (Jian et al., 2000)’s instrument, “The system’s actions will have a harmful or injurious outcome.” In our context, the outcome for the user is their earnings from driving while following the recommended schedule. We reworded the question to assess the outcome retrospectively.
9.
If there were any, please identify some of the driving times recommended by the tool that did not align with your expectations.
When was the time?
In 1 or 2 sentences, why did it not align with your expectations?
10.
Do you have any other questions or comments regarding this tool that you would like to share with us?
Appendix B Interview Scripts
B.1. Pilot
B.1.1. Formative Questions
Since the interviews were semi-structured, the script below focuses on the guiding questions that we asked participants. The interviewer also probed participants further depending on their responses.
Please tell us what you like most about driving for ridesharing and/or delivery services.
Please tell us what you like least about driving for ridesharing and/or delivery services.
When you drive, how important is it to you that you make some minimum amount of money daily/weekly?
When you drive, do you have an accurate sense of how much money you will make?
Do you try to stick to a regular routine for times and places to drive?
Are you happy with your current routines in terms of when and where you drive?
Would getting recommendations for times to drive would be helpful to you?
B.1.2. Evaluative Questions
We will show you a tool that can suggest personalised driving schedules. The tool will ask you some questions about your availability and preferences, as well as revenue targets that you might have. Different people might want to use the tool differently. However, we expect a typical user to use it as follows.
First, they would fill in some information about when they are available during the week, along with either how long they want to work or how much they want to make. The tool will then suggest a recommended schedule for the entire week. As they return to the tool every day to plan out their schedules, users will have the opportunity to interact with the tool, tweaking their availability and possible revenue targets to see how the recommendations change.
The interviewer begins screensharing the constraint page prototype.
Here’s the initial page of the tool that lets you specify your availability and goals.
Do you believe you understand what is being shown on this page?
Do you feel that this tool is asking you the right questions about your availability and goals?
Are there other important questions that you wish the tool would ask?
Think about your upcoming week. Using this screen, please tell us what information you think you would want to enter to get a useful recommendation. We will click on the page for you.
The interviewer switches to the schedule page prototype.
Here is an example of a recommended schedule that the tool would generate based on the information you just provided.
Do you feel that you understand what the recommended schedule is suggesting?
What part of the recommended schedule do you feel is the most useful?
What part of the recommended schedule do you feel is the least useful? Is there anything that’s missing from the schedule?
Do you feel that the recommended schedule gives you enough information to decide whether you would want to follow it?
Finally, we’d like to ask about your overall opinion of the tool.
What did you like about this tool?
What did you dislike about this tool?
Did you feel that interacting with the tool took too much time, or that it was too complicated or confusing for you? Why or why not?
What sort of information would increase the chance that you want to use this tool and follow its recommendations? How big of a difference do you think that having this information would make?
Do you believe that drivers would generally find a tool like this to be useful for when they’re planning their driving? Why or why not?
B.2. Longitudinal Trials
The script below focuses on the guiding questions that we asked participants. Some typical probing questions are also listed as sub-bullets.
B.2.1. Formative Questions
Let’s start with talking about your driving for rideshare/delivery services in general.
Could you start by telling us why you are driving?
Is it primary or supplemental income?
What other commitments do you balance it with (jobs, family, hobbies)?
To what extent do you rely on making a target amount when you are driving?
Can you talk through your typical process for deciding when to drive?
B.2.2. Feedback on Constraint Design
Now, let’s think back to the times when you were interacting with the tool, particularly when it asked you to enter your availability and goals.
How similar or different were the tool’s questions to the way you typically make these decisions?
Did you feel like you were able to use the tool to adequately specify your main considerations for when you’d like to drive?
Were you ever unsure of what information the tool was asking for?
Would you have preferred the tool to ask for information differently, or to ask for different information?
Did you feel like you were able to influence the recommended schedule that the tool generated for you?
Did you try to experiment with entering in different information?
Did you feel that interacting with the tool took too much time, or that it was too complicated or confusing for you?
Could you see yourself spending more time interacting with the tool than you did (e.g., to enter more details)? Why or why not?
B.2.3. Feedback on Schedules
Now, let’s talk about your how the tool’s recommended schedules may or may not have influenced your driving activity over the last few days.
Did you find that the recommended schedules made sense?
Did you feel that you understood how the tool used your answers to generate schedules? Why or why not?
To what extent did you rely on the email reminders of the schedules?
Did you ever miss the email reminders?
When did you typically check the schedule, if at all?
If you saw the recommended schedules, how did they impact your process for deciding when to work?
To what extent did you follow the schedules?
Were there times at which you prioritised your own intuition over the schedules?
If so, were there times at which you wished you followed the schedule more closely? Why or why not?
How did your response to the recommended schedules change throughout the week, if at all?
Did you look at the schedules more or less as time went on?
Were there any particular days on which you wanted to check the schedule more? Why or why not?
Did you feel that the tool gave you more or different information than you would otherwise get from the services that you drive for/from Gridwise? Why or why not?
Are there some additional details which could have increased the chance that you followed the recommended schedules?
For example, would you have preferred to see the schedule for the entire week on every day?
Did the recommended schedules lead you to drive at different times and/or locations than before?
Did this happen early on or later?
At what times of day?
When you followed the recommended schedules, did you feel that you ended up making more money, less money, or about the same relative to before?
How closely do you track your earnings in general?
Did you track your earnings more closely when using the tool?
B.2.4. Feedback on Estimates
How much did you focus on the tool’s estimates for how much you could earn?
Did you feel like you could rely on the estimates to achieve your earning goals?
Did you feel that these estimates were meant to be accurate projections of how much you could earn, or that they were rough ballpark figures?
In general, did you feel that the estimated earnings had the right level of detail, or would you have liked to see additional information?
If participants were in Conditions 1 or 2 Would you have preferred to see a range for how much you could earn?
If participants were in Conditions 3 or 4 Would you have preferred to see a single number for how much you could earn?
The interviewer selects a particular day on which the participant interacted with the tool. If earnings data was available, this was a day on which the participant earned more than the tool’s estimate; otherwise, this was the sixth day of their interaction with the tool.
Let’s talk about [weekday], Day [day] of your interaction with the tool.
Do you recall the extent to which you looked at the recommended schedule?
If you did, do you remember how you decided whether you wanted to follow it? Was this influenced by how much the tool estimated your earnings to be? Why or why not? What did you think of the estimate that the tool gave you?
If you didn’t, do you happen to remember why? Was this influenced by how much you earned on the previous day? Why or why not?
Do you recall whether you made more or less than the tool estimated on that day?
If there were any differences, do you have any idea why?
Did that influence your decision to look at the recommended schedule for the next day? Why or why not?
We checked your records briefly and found that you earned $xxx.xx, compared to the tool’s estimate of $xxx.xx. Does that change how you feel at all?
B.2.5. Overall Thoughts
Now, we’d like to wrap up with a few general questions about the tool.
Did you feel that the time you spent interacting with the tool was worthwhile or not worthwhile? Why or why not?
If you had the option of using a tool like this one, what are the chances that you might actually use it to decide your driving schedule in the future? Why or why not?
Beyond what you’ve mentioned already, is there anything else you believe might increase the chance that you would use this tool in the future?
Do you have any other questions, comments, or concerns?
Appendix C Confounding Variable Analysis
In this section, we detail the models we used to analyse whether several confounding variables could have impacted our measurements. We use the variable notation introduced in Section 6.2.1; for a participant’s selected constraint variant; for the participant’s city; and for the participant’s level of motivation, on a scale of 1–5.
C.1. Constraint Variants
To ascertain that trust and reliance measures did not significantly vary between different variants of constraints (maximum hours or minimum earnings/daily or weekly), we encoded these variants of the constraints as a categorical variable , and added it as linear terms to the best-fitting models for each measurement:
Survey 2 Trust (Table S1): OLS model from RQ3
Survey 3 Trust (Table S2): LMM from RQ1
Survey 3 Reliance (Table S3): LMM modified from RQ2
Survey 4 Trust (Table S4): OLS model from RQ1
These revised models generally led to worse AIC values (with models including interactions of being worse still), and most of the coefficients for the constraint variants were not significant. As noted in Section 6.2.1, one exception was the max/day group, who exhibited significantly higher reliance for Survey 3. The AIC was slightly improved by adding the constraints to this model. Fitted coefficients for each model are shown in the tables listed above.
Coefficient | |||
Intercept | 3.1846 | 0.376 | 0.000*** |
0.4304 | 0.423 | 0.315 | |
0.0459 | 0.338 | 0.893 | |
0.0317 | 0.433 | 0.942 | |
0.3341 | 0.383 | 0.387 | |
0.9385 | 0.403 | 0.024* | |
0.1244 | 0.409 | 0.762 |
Table of ordinary least squares model coefficients for pre-interaction Survey 2 trust, with constraint added. Significant coefficients are the intercept (3.1846) and Condition 3 (0.9385), but not .
Coefficient | |||
Fixed effects | |||
Intercept | 2.992 | 0.271 | 0.000*** |
0.216 | 0.268 | 0.421 | |
0.303 | 0.359 | 0.399 | |
0.540 | 0.403 | 0.180 | |
0.129 | 0.044 | 0.003** | |
0.005 | 0.002 | 0.003** | |
-0.001 | 0.000 | 0.008** | |
0.517 | 0.232 | 0.026* | |
-0.116 | 0.050 | 0.019* | |
Random effect | 0.499 | 0.246 |
Table of linear mixed model coefficients (intercept, fixed effects, and random effects) for daily Survey 3 trust, with constraint added. Significant coefficients are the intercept (2.992), (0.129), (0.005), the interaction of and (-0.001), (0.517), and the interaction of and (-0.116), but not .
Coefficient | |||
Fixed effects | |||
Intercept | 0.431 | 0.088 | 0.000*** |
0.133 | 0.101 | 0.186 | |
0.322 | 0.134 | 0.016* | |
-0.144 | 0.153 | 0.346 | |
0.230 | 0.073 | 0.002** | |
0.058 | 0.066 | 0.383 | |
Random effect | 0.059 | 0.073 |
Table of linear mixed model coefficients (intercept, fixed effects, and random effects) for daily Survey 3 reliance, with constraint added. Significant coefficients are the intercept (0.431), (0.23), and = max/day (0.322).
Coefficient | |||
Intercept | 3.539 | 0.345 | 0.000*** |
-0.021 | 0.256 | 0.934 | |
0.267 | 0.323 | 0.416 | |
-1.139 | 0.732 | 0.131 | |
0.274 | 0.349 | 0.439 |
Table of ordinary least squares model coefficients for post-interaction Survey 4 trust, with constraint added. The intercept (3.539) is significant, but the coefficients for (0.274) and are not.
C.2. Cities
To ascertain that trust and reliance measures did not significantly vary based on the city in which participants were driving (Los Angeles, New York, Chicago, or Houston), we similarly added a categorical variable for the city to the best-fitting models:
Survey 2 Trust (Table S5): OLS model
Survey 3 Trust (Table S6): LMM
Survey 3 Reliance (Table S7): LMM
Survey 4 Trust (Table S8): OLS model
These revised models all led to worse AIC values, and none of the coefficients for the cities were significant.
Coefficient | |||
Intercept | 3.0656 | 0.467 | 0.000*** |
0.4695 | 0.497 | 0.350 | |
0.2597 | 0.417 | 0.537 | |
0.3459 | 0.494 | 0.488 | |
0.2646 | 0.394 | 0.505 | |
0.8850 | 0.430 | 0.045* | |
0.0353 | 0.405 | 0.931 |
Table of ordinary least squares model coefficients for pre-interaction Survey 2 trust, with city added. Significant coefficients are the intercept (3.0656) and Condition 3 (0.885), but not .
Coefficient | |||
Fixed effects | |||
Intercept | 3.121 | 0.371 | 0.000*** |
0.174 | 0.440 | 0.693 | |
0.015 | 0.351 | 0.966 | |
0.053 | 0.402 | 0.895 | |
0.132 | 0.044 | 0.003** | |
0.004 | 0.002 | 0.004** | |
-0.001 | 0.000 | 0.011* | |
0.533 | 0.233 | 0.022* | |
-0.120 | 0.050 | 0.016* | |
Random effect | 0.525 | 0.257 |
Table of linear mixed model coefficients (intercept, fixed effects, and random effects) for daily Survey 3 trust, with city added. Significant coefficients are the intercept (3.121), (0.132), (0.004), the interaction of and (-0.001), (0.533), and the interaction of and (-0.12), but not .
Coefficient | |||
Fixed effects | |||
Intercept | 0.539 | 0.137 | 0.000*** |
0.002 | 0.179 | 0.990 | |
-0.033 | 0.143 | 0.818 | |
-0.023 | 0.163 | 0.888 | |
0.213 | 0.073 | 0.004** | |
0.067 | 0.066 | 0.308 | |
Random effect | 0.077 | 0.090 |
Table of linear mixed model coefficients (intercept, fixed effects, and random effects) for daily Survey 3 reliance, with city added. Significant coefficients are the intercept (0.539) and (0.213), but not .
Coefficient | |||
Intercept | 3.4142 | 0.389 | 0.000*** |
0.1228 | 0.410 | 0.767 | |
0.0096 | 0.321 | 0.976 | |
-0.3000 | 0.388 | 0.445 | |
0.4629 | 0.331 | 0.173 |
Table of ordinary least squares model coefficients for post-interaction Survey 4 trust, with city added. The intercept (3.4142) is significant, but the coefficients for (0.4629) and are not.
C.3. Motivation
To ascertain that trust and reliance measures did not significantly vary with participants’ level of motivation, we needed to first select a survey item that would allow us to assess this subjective dimension independent of the participants’ perception of the tool. For this reason, we selected Question 5 of the formative Survey 1 (Section A.1), “When I drive, it’s important to me that I make some minimum amount of money.” We encoded the Likert responses to this question as an interval variable and added it as linear terms to the best-fitting models for each measurement:
Survey 2 Trust (Table S9): OLS model from RQ3
Survey 3 Trust (Table S10): LMM from RQ1
Survey 3 Reliance (Table S11): LMM modified from RQ2
Survey 4 Trust (Table S12): OLS model from RQ1
These revised models again generally led to worse AIC values, and most of the coefficients for the motivation variable were not significant. As noted in Section 6.2.1, one exception was the pre-interaction trust score, which had a significant positive correlation with the motivation variable. The AIC was improved by adding the motivation variable to this model. However, no such effects were evident for the daily and post-interaction measures. This suggests that the initial effects of the participants’ motivation on trust did not persist over repeated interactions. Fitted coefficients for each model are shown in the tables listed above.
Coefficient | |||
Intercept | 2.1844 | 0.475 | 0.000*** |
0.3001 | 0.105 | 0.007** | |
0.0725 | 0.352 | 0.838 | |
0.7487 | 0.359 | 0.043* | |
-0.0430 | 0.358 | 0.905 |
Table of ordinary least squares model coefficients for pre-interaction Survey 2 trust, with motivation added. Significant coefficients are the intercept (2.1844), Condition 3 (0.7487), and (0.3001).
Coefficient | |||
Fixed effects | |||
Intercept | 2.990 | 0.570 | 0.000*** |
0.046 | 0.119 | 0.699 | |
0.126 | 0.044 | 0.004** | |
0.002 | 0.001 | 0.001*** | |
-0.000 | 0.000 | 0.004** | |
0.530 | 0.230 | 0.021* | |
-0.116 | 0.049 | 0.018* | |
Random effect | 0.530 | 0.261 |
Table of linear mixed model coefficients (intercept, fixed effects, and random effects) for daily Survey 3 trust, with motivation added. Significant coefficients are the intercept (2.99), (0.126), (0.002), the interaction of and (-0.000), (0.53), and the interaction of and (-0.116), but not .
Coefficient | |||
Fixed effects | |||
Intercept | 0.383 | 0.234 | 0.102 |
0.030 | 0.050 | 0.550 | |
0.211 | 0.071 | 0.003** | |
0.074 | 0.064 | 0.250 | |
Random effect | 0.081 | 0.094 |
Table of linear mixed model coefficients (intercept, fixed effects, and random effects) for daily Survey 3 reliance, with motivation added. The fixed effect for (0.211) is significant, but not .
Coefficient | |||
Intercept | 3.9645 | 0.676 | 0.000*** |
-0.1259 | 0.127 | 0.330 | |
0.4258 | 0.317 | 0.189 |
Table of ordinary least squares model coefficients for post-interaction Survey 4 trust, with motivation added. The intercept (3.9645) is significant, but the coefficients for (0.4258) and are not.
Appendix D Final Tool Design
The following figures show screenshots of the web-based tool used for the longitudinal trials (Section 6), showing the constraints and recommended schedule of interview participant F1.
Screenshot of a webpage titled “Tool Interaction”. Below a set of instructions, the constraint page is shown in a grey iframe. Various inputs for constraints are set to values entered by interview participant F1. In response to “Would you like to specify a maximum number of hours that you would like to drive, or would you prefer to specify a minimum amount of money that you would like to earn?”, F1 selected the latter. In response to “Would you like to set these constraints for each day individually, or for the entire week?”, F1 selected the latter. In response to “What is the minimum amount of money you would like to earn for the entire week?”, F1 entered $1 400. In response to “When are you available during the week?”, F1 selected time slots (as indicated by blue cells in a tabular schedule) corresponding to 3 am to 7 am, 8 am to 4 pm, and 10 pm to 12 am on every day of the week.
D.1. Schedule Page Conditions
Screenshot of a webpage titled “Tool Interaction”. Below a set of instructions, the schedule page in Condition 1 for interview participant F1 is shown in a grey iframe. The top of the schedule page reads, “Based on historical data, it is estimated that you will earn a weekly average of”, followed by an estimate of $1 412 below. Further below, a tabular schedule is shown for each day of the week. Each cell reads, “Average”, followed by the estimated hourly earning in bold. Recommended time slots are indicated by blue cells, non-recommended time slots during which F1 is still available are indicated by dark text, and time slots during which F1 is not available are indicated by light text.
Screenshot of a webpage titled “Tool Interaction”. Below a set of instructions, the schedule page in Condition 2 for interview participant F1 is shown in a grey iframe. The top of the schedule page reads, “Based on historical data, it is estimated that you will earn”. Unlike Condition 1, this is followed by a table that reads, “An average of”, alongside estimates for each day: $210 for Monday, $194 for Tuesday, $175 for Wednesday, $173 for Thursday, $194 for Friday, $226 for Saturday, and $237 for Sunday. Further below, a tabular schedule is shown for each day of the week. Each cell reads, “Average”, followed by the estimated hourly earning in bold. Recommended time slots are indicated by blue cells, non-recommended time slots during which F1 is still available are indicated by dark text, and time slots during which F1 is not available are indicated by light text.
Screenshot of a webpage titled “Tool Interaction”. Below a set of instructions, the schedule page in Condition 3 for interview participant F1 is shown in a grey iframe. The top of the schedule page reads, “Based on historical data, it is estimated that you will earn”. Unlike Conditions 1 and 2, this is followed by a table that reads, “On an average week: $1 412”, “On a bad week: $391”, and “On a good week: $3 032”. Further below, a tabular schedule is shown for each day of the week. Each cell reads, “Average”, followed by the estimated hourly earning in bold. Unlike Conditions 1 and 2, this is also followed by “Low”, with the 10th percentile of earnings, and “High”, with the 90th percentile of earnings. Recommended time slots are indicated by blue cells, non-recommended time slots during which F1 is still available are indicated by dark text, and time slots during which F1 is not available are indicated by light text.
Screenshot of a webpage titled “Tool Interaction”. Below a set of instructions, the schedule page in Condition 4 for interview participant F1 is shown in a grey iframe. Unlike Conditions 1, 2, and 3, the top of the schedule page reads, “In general, every driver may encounter different conditions while they are driving, which leads to differences in earnings from individual to individual and from one day to the next. On average, based on historical data, a driver following this schedule will earn”. Also unlike Conditions 1 and 2, this is followed by a table that reads, “On an average week: $1 412”, “On a bad week: $391”, and “On a good week: $3 032”. Further below, a tabular schedule is shown for each day of the week. Each cell reads, “Average”, followed by the estimated hourly earning in bold. Unlike Conditions 1 and 2, this is also followed by “Low”, with the 10th percentile of earnings, and “High”, with the 90th percentile of earnings. Recommended time slots are indicated by blue cells, non-recommended time slots during which F1 is still available are indicated by dark text, and time slots during which F1 is not available are indicated by light text.
Appendix E Figma Prototype Design
The following figures show screenshots of the Figma prototype used for the pilot interviews (Section 5).
Screenshot of a webpage titled “Shift Recommendation Tool”. The constraint page is shown, with various inputs being set to mock values. In response to “Would you like to set a maximum number of hours that you would like to drive to earn as much money as possible, or a minimum amount of money that you would like to earn in as few hours as possible?”, the latter is selected. In response to “Would you like to set these constraints for each day individually, or for the entire week?”, the former is selected. In response to “What is the minimum amount of money you would like to earn for each day?”, $140–$150 is entered for each day of the week. In response to “When are you available during the week?”, time slots corresponding to 7 am to 2 pm on every day of the week are selected (as indicated by dark cells in a tabular schedule).
Screenshot of a webpage titled “Shift Recommendation Tool”. The schedule page is shown with mock estimates. The top of the schedule page reads, “Based on historical data, it is estimated that you will earn a weekly average of”. This is followed by a table that reads, “On an average week: $593”, “On a bad week: $152”, and “On a good week: $1 135”. Further below, a tabular schedule is shown for each day of the week. Each cell reads, “Average”, followed by the estimated hourly earning in bold. Recommended time slots are indicated by dark cells, and non-recommended time slots are indicated by light cells.